### v2

# Almost Surely Safe Alignment of Large Language Models at Inference-Time

Abstract
Even highly capable large language models (LLMs) can produce biased or unsafe responses, and alignment techniques, such as RLHF, aimed at mitigating this issue, are expensive and prone to overfitting as they retrain the LLM. This paper introduces a novel inference-time alignment approach that ensures LLMs generate safe responses almost surely, i.e., with a probability approaching one. We achieve this by framing the safe generation of inference-time responses as a constrained Markov decision process within the LLM’s latent space. Crucially, we augment a safety state that tracks the evolution of safety constraints and enables us to demonstrate formal safety guarantees upon solving the MDP in the latent space. Building on this foundation, we propose InferenceGuard, a practical implementation that safely aligns LLMs without modifying the model weights. Empirically, we demonstrate InferenceGuard effectively balances safety and task performance, outperforming existing inference-time alignment methods in generating safe and aligned responses.
Contains potentially harmful examples.
1 Introduction
LLMs have demonstrated impressive capabilities across a diverse set of tasks, such as summarization (Koh et al., 2022; Stiennon et al., 2020), code generation (Gao et al., 2023; Chen et al., 2021), and embodied robotics (Mower et al., 2024; Kim et al., 2024). However, since those models are primarily trained on vast, unsupervised datasets, their generated responses can often be biased, inaccurate, or harmful (Deshpande et al., 2023; Ganguli et al., 2022; Weidinger et al., 2021; Gehman et al., 2020). As a result, LLMs require alignment with human values and intentions to ensure their outputs are free from controversial content.
The traditional LLM alignment approach involves fine-tuning the model on human-labeled preference data. For instance, works such as (Ouyang et al., 2022; Christiano et al., 2017) use the reinforcement learning from human feedback (RLHF) framework to fine-tune LLMs, constructing a reward model from human feedback and optimizing the model using standard reinforcement learning algorithms like PPO (Schulman et al., 2017). More recent approaches, such as (Tutnov et al., 2025; Yin et al., 2024; Rafailov et al., 2023), bypass reward learning and instead align pre-trained models directly with human preferences.
RLHF alignment can be costly and risks overfitting partly since they modify the LLM’s model weights. To tackle these challenges, alternative approaches adjust the model’s responses directly at inference time while leaving the pre-trained model weights fixed. Several techniques have been proposed for this purpose, such as Best-of-N (Nakano et al., 2021; Stiennon et al., 2020; Touvron et al., 2023), FUDGE (Yang & Klein, 2021), COLD (Qin et al., 2022), CD-Q Mudgal et al. (2023), RE-control Kong et al. (2024), among others. Importantly, those techniques are designed modularly, allowing the alignment module to integrate seamlessly with the pre-trained model. This modularity enables flexible inference-time reconfigurability and quick adaptation to new reward models and datasets. Moreover, it reduces reliance on resource-intensive and often hard-to-stabilize RL processes inherent in the RLHF paradigm.
Despite the successes of inference-time alignment, these methods’ safety aspects have received limited attention so far. While some works have attempted to tackle those issues in inference-time-generated responses, they mainly focus on prompt-based alignment  (Hua et al., 2024; Zhang et al., 2024b), trainable safety classifiers  Niu et al. (2024); Zeng et al. (2024) or protections against adversarial attacks and jailbreaks  (Dong et al., 2024; Guo et al., 2024; Inan et al., 2023). That said, prompt-based methods cannot be guaranteed to consistently produce safe responses, as ensuring safety is heavily reliant on user intervention, requiring extensive engineering and expertise to manage the model’s output effectively. Trainable classifiers focus only on the safety of decoded responses using hidden states or virtual tokens, ignoring task alignment and lacking theoretical guarantees. Moreover, while adversarial robustness is crucial, our work focuses on the key challenge of generating inherently safe responses from the LLM.
In this work, we aim to develop LLMs that generate safe responses at inference time, following a principled approach that guarantees safety almost surely, i.e., with a probability approaching one. To do so, we reformulate the safe generation of inference-time responses as an instance of constrained Markov decision processes (cMDP). We map the cMDP to an unconstrained one through safety state augmentation, bypassing Lagrangian approaches’ limitations that struggle to balance reward maximization and safety feasibility. Focusing on a practical test-time inference algorithm, we adopt a critic-based approach to solve the augmented MDP, eliminating the need for gradients in the LLM. To ensure efficiency, we train our critic in the latent space of the LLM, keeping it small in size and fast during inference. This shift to the latent space complicates the theoretical framework, requiring extensions from (Hernández-Lerma & Muñoz de Ozak, 1992; Sootla et al., 2022). By doing so, we establish, for the first time, that one can guarantee almost sure safety in the original token space.
To leverage this theoretical guarantee in practice, we build upon the augmented MDP framework and introduce two novel implementations for safe inference-time alignment: one that learns a compact critic in the latent space for cases where safety costs can only be queried after the generation of complete responses and another that leverages direct cost queries for efficient inference-time optimization. Finally, we integrate these components into a lookahead algorithm (e.g., Beam Search or Blockwise Decoding (Mudgal et al., 2023)) proposing InferenceGuard. Our experiments tested InferenceGuard starting from safety-aligned and unaligned models. Our results achieved high safety rates—91.04% on Alpaca-7B and 100% on Beaver-7B-v3. Notably, this was accomplished while maintaining a strong balance with rewards, setting new state-of-the-art.
2 Background
2.1 LLMs as Stochastic Dynamical Systems
LLMs can be viewed as stochastic dynamical systems, where the model’s behavior evolves probabilistically over time, governed by its internal parameters and the inputs it receives. In this perspective, each new token is generated based on the model’s evolving hidden state (Kong et al., 2024; Zimmer et al., 2024). Formally, an LLM transitions as follows: $\left[\mathbf{h}_{t+1},\mathbf{o}_{t+1}\right]^{\mathsf{T}}=f_{\text{LLM}}(\mathbf{h}_{t},\text{y}_{t})$, with $\text{y}_{t}\sim\text{SoftMax}(\mathbf{W}\mathbf{o}_{t})$. Here, $f_{\text{LLM}}(\cdot)$, denotes the aggregation of all decoding layers, $\text{y}_{t}$ a generated token at each time step $t$, and $\mathbf{o}_{t}$ the logits which are linearly mapped by $\mathbf{W}$ to produce a probability distribution over the vocabulary space. Moreover, $\mathbf{h}_{t}$ comprises all key-value pairs accumulated from previous time steps  [^1] Note, $\mathbf{h}_{t}=\left\{\mathbf{K}^{(l)}_{j},\mathbf{V}_{j}^{(l)}\right\}_{l=1}^{L}$ for $j:[1:t]$, i.e., keys and values from all layers accumulated up to the previous time step.. The system evolves until the end-of-sequence (EOS) token is reached.
2.2 Test-Time Alignment of LLMs
Test-time alignment ensures that a pre-trained LLM generates outputs consistent with desired behaviors by solving a MDP, whose initial state is determined by the test prompt. Rewards/costs for alignment can come from various sources, including but not limited to human feedback (Tutnov et al., 2025; Zhong et al., 2024a), environmental feedback from the task or verifiers (Zeng et al., 2023; Yang et al., 2024; Trinh et al., 2024; An et al., 2024; Liang et al., 2024; Mower et al., 2024), or pre-trained reward models (Wang et al., 2024b; Zhang et al., 2024a; Li et al., 2025b).
Several approaches have been developed to address the challenge of test-time alignment. For instance, beam search with rewards (Choo et al., 2022) extends traditional beam search technique by integrating a reward signal to guide LLM decoding at test time. Monte Carlo tree search, on the other hand, takes a more exploratory approach by simulating potential future token sequences to find the path that maximizes the reward function (Zhang et al., 2024a). Best-of-N (BoN) generates multiple candidate sequences and selects the one with the highest reward  (Stiennon et al., 2020; Nakano et al., 2021; Touvron et al., 2023).
We focus on beam search and look-ahead methods for more scalability when developing InferenceGuard.
Figure 1: To aid readability, this figure summarizes the key transitions and notations used throughout the paper.
3 Safe Test-Time Alignment of LLMs
We frame the problem of safely aligning LLMs at test time as a constrained Markov decision process (cMDP). As noted in Achiam et al. (2017); Sootla et al. (2022) a cMDP is defined as the following tuple: $\mathcal{M}=\left\langle\mathcal{S},\mathcal{A},\mathcal{C}_{\text{task}},\mathcal{C}_{\text{safety}},\mathcal{P},\gamma\right\rangle$, with $\mathcal{S}$ and $\mathcal{A}$ denoting the state and action spaces, respectively. The cost function $\mathcal{C}_{\text{task}}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ dictates the task’s cost [^2] We define costs as negative rewards, which transforms the problem into an equivalent cost-based MDP., while $\mathcal{C}_{\text{safety}}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ represents a safety cost, which encodes the constraints that the actor must satisfy during inference. The transition model $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ captures the probability of transitioning to a new state given the current state and action. Meanwhile, the discount factor $\gamma\in[0,1)$ trades off immediate versus long-term rewards. The goal of constrained MDPs is to find a policy $\pi:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ that minimizes the task’s cost while simultaneously satisfying the safety constraints. Given a safety budget d, we write:
$\min_{\pi}$ $\mathbb{E}_{\mathcal{P},\pi}\Big{[}\sum_{t}\gamma^{t}\mathcal{C}_{\text{task}}(\mathbf{s}_{t},\mathbf{a}_{t})\Big{]}$ (1)
s.t. $\mathbb{E}_{\mathcal{P},\pi}\Big{[}\sum_{t}\gamma^{t}\mathcal{C}_{\text{safety}}(\mathbf{s}_{t},\mathbf{a}_{t})\Big{]}\leq d,$
3.1 Safe Test-Time Alignment as cMDPs
We treat the generation process of safe test-time alignment of LLMs as the solution to a specific cMDP. We introduce our state variable $\mathbf{s}_{t}=\{\mathbf{x},\mathbf{y}_{<t}\}$, which combines the input prompt $\mathbf{x}$ with the tokens (or partial responses) decoded until step $t$. Our policy generates a new token $\text{y}_{t}$ that we treat as an action in the model’s decision-making process. The transition function $\mathcal{P}$ of our MDP is deterministic, where the state $\mathbf{s}_{t}$ is updated by incorporating the generated action $\text{y}_{t}$, i.e., $\mathbf{s}_{t+1}=\mathbf{s}_{t}\oplus\text{y}_{t}=\{\mathbf{x},\mathbf{y}_{\leq t}\}$. We also assume the existence of two cost functions $\mathcal{C}_{\text{task}}$ and $\mathcal{C}_{\text{safety}}$ to assess the correctness and safety of the LLM’s responses. As described in Section 2.2, those functions can originate from various sources, such as human feedback, environmental verifiers, or pre-trained models. While we conduct experiments with these functions being LLMs (see Section 5), our method can be equally applied across various types of task and safety signals.
Regarding the task’s cost, we assume the availability of a function $\mathcal{C}_{\text{task}}$ that evaluates the alignment of the LLM with the given task. This function assigns costs to the partial response based on the input prompt $\mathbf{x}$ such that:
$\mathcal{C}_{\text{task}}([\mathbf{x},\mathbf{y}_{\leq t}]):=\begin{cases}0&\text{if }\text{y}_{t}\neq\texttt{EOS}\\ c_{\text{task}}([\mathbf{x},\mathbf{y}_{\leq t}])&\text{if }\text{y}_{t}=\texttt{EOS}\end{cases}$ (2)
For the safety cost $\mathcal{C}_{\text{safety}}$, we assume the function assigns non-zero costs to any partial answer without waiting for the final token. This is crucial because we want to flag unsafe responses early rather than waiting until the end of the generation process. Many pre-trained models are available for this purpose on Hugging Face, which we can leverage—more details can be found in Section 5. With this, we write safe test-time alignment as an instance of Equation 1:
$\min_{\pi}$ $\mathbb{E}_{\pi}\Big{[}\sum_{t}\gamma^{t}\mathcal{C}_{\text{task}}(\overbrace{\{\mathbf{x},\mathbf{y}_{<t}\}}^{\mathbf{s}_{t}},\text{y}_{t})\Big{]}$ (3)
s.t. $\mathbb{E}_{\pi}\Big{[}\sum_{t}\gamma^{t}\mathcal{C}_{\text{safe}}(\{\mathbf{x},\mathbf{y}_{<t}\},\text{y}_{t})\Big{]}\leq d.$
The above objective aims to minimize the task cost $\mathcal{C}_{\text{task}}$ while ensuring that the safety cost does not exceed a predefined budget $d$. The expectation is taken over the actions (tokens) generated at each timestep.
3.2 State Augmented Safe Inference-Time Alignment
We could technically use off-the-shelf algorithms to solve Equation 3, such as applying a Lagrangian approach as proposed in Dai et al. (2023). However, there are two main issues with using these standard algorithms. First, they generally require gradients in the model itself—specifically, the LLM—which we want to avoid since our goal is to perform inference-time alignment without retraining the model. Second, these methods rely on a tunable Lagrangian multiplier, making it challenging to maximize rewards while satisfying almost sure constraints optimally.
Instead of using a Lagrangian approach, we take a different direction by augmenting the state space and extending the method proposed by (Sootla et al., 2022) to large language models. In our approach, we augment the state space of the constrained MDP with a “constraint tracker”, effectively transforming the problem into an unconstrained one. This allows us to apply Bellman equations and conduct rigorous proofs with almost sure constraint satisfaction results. However, applying the techniques and proofs from (Sootla et al., 2022) to our test-time setting is not entirely straightforward due to two main challenges: first, the differences in the constrained MDP setting, and second, the process by which we train critics, as we will demonstrate next.
Augmenting the State Space. The following exposition builds on (Sootla et al., 2022), extending their method to address LLM-specific challengers, an area they did not cover. The core idea is to transform the constrained MDP into an unconstrained one by augmenting the state with an additional variable that tracks the remaining budget of the constraint. While doing so, we must ensure that: PI) our augmented state variable tracks the constraints and maintains the Markovian nature of transition dynamics; and PII) our task cost $\mathcal{C}_{\text{task}}$ accounts for this new state representation and is correctly computed w.r.t. the augmented state.
To solve PI, we investigate the evolution of the constraint in Equation 3 and track a scaled-version of the remaining safety budget $\mathbf{\omega}_{t}=d-\sum_{k=1}^{t}\gamma^{k}\mathcal{C}_{\text{safety}}(\{\mathbf{x},\mathbf{y}_{<k}\},\text{y}_{k})$ that is defined as $\text{z}_{t}=\mathbf{\omega}_{t-1}/\gamma^{t}$. The update of $\text{z}_{t}$ satisfies:
$\text{z}_{t+1}$ $=(\mathbf{\omega}_{t-1}-\gamma^{t}\mathcal{C}_{\text{safety}}(\{\mathbf{x},\mathbf{y}_{<t}\},\text{y}_{t}))/\gamma^{t+1}$ (4)
$=(\text{z}_{t}-\mathcal{C}_{\text{safety}}(\{\mathbf{x},\mathbf{y}_{<t}\},\text{y}_{t}))/\gamma,\ \ \text{with $\text{z}_{0}=d$.}$
Since the dynamics of $\text{z}_{t}$ are Markovian due to the dependence of $\text{z}_{t+1}$ only on $\text{z}_{t}$, $\text{y}_{t}$ and current the state $\{\mathbf{x},\mathbf{y}_{<t}\}$, we can easily augment our original state space with $\text{z}_{t}$, such that $\tilde{\mathbf{s}}_{t}=[\mathbf{s}_{t},\text{z}_{t}]=[\{\mathbf{x},\mathbf{y}_{<t}\},\text{z}_{t}]$. The original dynamics can also be redefined to accommodate for $\tilde{\mathbf{s}}_{t}$:
[…]
Concerning PII, we note that enforcing the original constraint in Equation 3 is equivalent to enforcing an infinite number of the following constraints:
$\sum_{k=0}^{t}\gamma^{k}\mathcal{C}_{\text{safety}}(\{\mathbf{x},\mathbf{y}_{<k}\},\text{y}_{k})\leq d\ \ \forall t\geq 1.$ (5)
As noted in (Sootla et al., 2022), this observation holds when the instantaneous costs are nonnegative, ensuring that the accumulated safety cost cannot decrease. In our case, it is natural to assume that the costs are nonnegative for LLMs, as safety violations or misalignments in the output typically incur a penalty, reflecting the negative impact on the model’s performance or ethical standards.
Clearly, if we enforce $\text{z}_{t}\geq 0$ for all $t\geq 0$, we automatically get that $\mathbf{\omega}_{t}=d-\sum_{k=0}^{t}\gamma^{k}\mathcal{C}_{\text{safety}}(\{\mathbf{x},\mathbf{y}_{<k}\},\text{y}_{k})\geq 0$ for all $t\geq 0$, thus satisfying the infinite constraints in Equation 5. We can do so by reshaping the tasks’s instantaneous cost to account for the safety constraints:
$\tilde{\mathcal{C}}_{\text{task}}^{\infty}(\tilde{\mathbf{s}}_{t},\text{y}_{t}):=\begin{cases}\mathcal{C}_{\text{task}}([\mathbf{x},\mathbf{y_{\leq t}}])&\text{z}_{t}>0\\ +\infty&\ \text{z}_{t}\leq 0,\end{cases}$ (6)
[…]
$\min_{\pi}\mathbb{E}_{\pi}\Big{[}\sum_{t}\gamma^{t}\tilde{\mathcal{C}}_{\text{task}}^{\infty}(\tilde{\mathbf{s}}_{t},\text{y}_{t})\Big{]}.$ (7)
Using gradient-based techniques, one could optimize the augmented MDP in Equation 7. However, since our goal is to enable safety at test time without retraining, we adopt a critic-based approach that does not require gradients during inference, as we show next.
4 InferenceGuard: Safety at Test-Time
When designing our critic, we considered several crucial factors for test-time inference. These included its size, ease of training for quick adaptation, and flexibility to operate in real-time without significant latency. As such, we chose to train the critic in the latent space of the LLM rather than directly in the textual space, enabling a more efficient solution that meets the constraints of test-time alignment.
[…]
To fully define our critic, we require a representation of the embedding of our augmented state $\tilde{\mathbf{s}}_{t}=[\mathbf{s}_{t},\text{z}_{t}]$ within the latent space. As noted above, we can acquire $(\mathbf{h}_{t},\mathbf{o}_{t})$ from the transformer architecture. We call this mapping $\phi$, whereby $(\mathbf{h}_{t},\mathbf{o}_{t})=\phi(\{\mathbf{x},\mathbf{y}_{<t}\})$. Furthermore, we use an identity mapping to embed $\text{z}_{t}$, which enables us to input the actual tracking of the constraints directly to the critic without any loss of information.
4.1 Theoretical Insights
In this section, we show that optimizing in the latent space preserves safety constraints in the original token space, and we prove that our approach guarantees almost sure safety.
We should answer two essential questions: i) Can we still compute an optimal policy in the latent space? and ii) If we enforce safety constraints in the latent space, do they still hold in the original token space? While the prior work in (Sootla et al., 2022) established theoretical results for safety-augmented MDPs in standard (non-LLM) RL settings, their work does not address how guarantees in the latent space translate to the original token space. To handle those problems, we extend the theorems from Hernández-Lerma & Muñoz de Ozak (1992); Sootla et al. (2022) to ensure the following three properties:
  • Prop I) The latent MDP indeed satisfies the Bellman equations (Theorem 4.1 (a)) and, hence, allows us to compute an optimal policy in this space,
  • Prop II) Policies and value functions in the latent space are valid in the original token space, implying that optimizing in the latent space preserves the constraints in original token space (Theorem 4.1 (b,c))
  • Prop III) The resulting policy satisfies safety constraints almost surely (Theorem 4.1), meaning if a policy is safe in the latent and original token space with finite expected cost w.r.t. Equation 7, it is also almost surely safe in the actual LLM token space.
We begin by defining the latent space MDP’s cost and transition function:
Definition 4.1.
$\exists\phi(\cdot)$ and functions $\bar{\mathcal{C}}^{n}_{\text{task}}$ and $\bar{\mathcal{P}}$ such that:
[…]
$\bar{\mathcal{P}}(\phi(\{\mathbf{x},\mathbf{y}_{\leq t}\}),\text{z}_{t+1}|\phi(\{\mathbf{x},\mathbf{y}_{<t}\}),\text{z}_{t},\text{y}_{t})={\mathcal{P}}(\tilde{\mathbf{s}}_{t+1}|\tilde{\mathbf{s}}_{t},\text{y}_{t}),$
where $\tilde{\mathbf{s}}_{t}$ is the augmented state in the original token space.
According to Definition 4.1, we ensure that the cost incurred by the augmented state $\tilde{\mathbf{s}}_{t}=[\{\mathbf{x},\mathbf{y}_{<t}\},\text{z}_{t}]$ w.r.t. $\tilde{\mathcal{C}}^{n}_{\text{task}}$ is equal to the latent cost incurred by the latent state $[\phi(\{\mathbf{x},\mathbf{y}_{<t}\}),\text{z}_{t}]$ w.r.t $\bar{\mathcal{C}}_{\text{task}}^{n}$. Additionally, we ensure that the transition dynamics of the augmented state in the original token space and the corresponding latent state in the latent space are equivalent.
This equivalence enables us to derive an optimal policy for the latent MDP and apply it to minimize the cost objective in the original augmented MDP (see Equation 7). We proceed to analyze the existence of such an optimal policy in the latent space through the following standard assumptions (Sootla et al., 2022) on $\bar{\mathcal{C}}^{n}_{\text{task}}$, and $\bar{\mathcal{P}}$: A1. The function $\bar{\mathcal{C}}_{\text{task}}^{n}(\mathbf{h},\mathbf{o},\text{z},\text{y})$ is bounded, measurable, nonnegative, lower semi-continuous w.r.t. $(\mathbf{h},\mathbf{o},\text{z})$ for a given y, and A2. The transition law $\bar{{\mathcal{P}}}$ is weakly continuous for any y.
Next, we define $\bar{\pi}$ as a policy in the latent space that maps $(\mathbf{h},\mathbf{o},\text{z})\rightarrow\text{y}$, and its value function for an initial state $(\mathbf{h}_{0},\mathbf{o}_{0},\text{z}_{0})$ as follows: $\bar{V}^{n}(\bar{\pi},\mathbf{h}_{0},\mathbf{o}_{0},\text{z}_{0})=\mathbb{E}_{\bar{\pi}}\Big{[}\sum_{t=0}^{\infty}\gamma^{t}\bar{\mathcal{C}}_{\text{task}}^{n}(\mathbf{h}_{t},\mathbf{o}_{t},\text{z},\text{y})\Big{]}$. Then, one can define the optimal value function:
$\bar{V}^{\star,n}(\mathbf{h},\mathbf{o},\text{z})=\min_{\bar{\pi}}\bar{V}^{n}(\bar{\pi},\mathbf{h},\mathbf{o},\text{z}).$ (8)
Since we cannot optimize directly in the original constrained MDP, we first show that solving for an optimal policy in the latent MDP preserves key properties of the original problem. The following theorem formalizes this by proving the existence of the optimal policy and its mapping to the original MDP. The proof is in Appendix B.3.
{restatable}
thmthmsauteequivalence(Optimality in the Latent Space) Given the latent MDP in Definition 4.1 and with A1-A2, we can show:
  1. a)
    (Prop I) For any finite $n$, the Bellman equation holds, i.e., there exists a function $\bar{V}^{\star,n}(\mathbf{h},\mathbf{o},\text{z})$ such that:
    $\bar{V}^{\star,n}(\mathbf{h},\mathbf{o},\text{z})=\min_{\text{y}\in{\mathcal{V}}}\Big{(}\bar{\mathcal{C}}^{n}_{\text{task}}(\mathbf{h},\mathbf{o},\text{z},\text{y})$
    $\hskip 130.0002pt+\gamma\bar{V}^{\star,n}(\mathbf{h}^{\prime},\mathbf{o}^{\prime},\text{z}^{\prime})\Big{)},$
    $\text{such that}\ \ (\mathbf{h}^{\prime},\mathbf{o}^{\prime},\text{z}^{\prime})\sim\bar{{\mathcal{P}}}(\cdot|\mathbf{h},\mathbf{o},\text{z},\text{y})$
    Furthermore, the optimal policy solving Equation 8 has the representation $y\sim\bar{\pi}^{\star,n}(\cdot\mid\mathbf{h},\mathbf{o},\text{z})$;
  2. b)
    (Prop II) The optimal value functions $\bar{V}^{\star,n}$ converge monotonically to $\bar{V}^{\star,\infty}$.
  3. c)
    (Prop II) The optimal policy in the latent space $\bar{\pi}^{\star,n}$ is also optimal in the original token space if used as $\bar{\pi}^{\star,n}(\phi(\cdot))$, minimizing Equation 7, even as $n\rightarrow\infty$.
The above theorem ensures that finding and securing the existence of the optimal policy in the latent space is sufficient to solve Equation 7 optimally [^5] As noted in Section 3.2, we analyze the transition from a large $n$ to $\infty$, and confirm that the results hold even for $\tilde{\mathcal{C}}_{\text{task}}^{\infty}$.. Informally, the latent space acts as a faithful representation, preserving constraints and making optimization computationally efficient. This implies that the optimal policies and value functions in the latent space remain valid in the original space.
Almost Sure Guarantee. Now, we derive Prop III that ensures the safety cost constraints are almost surely satisfied. This is more challenging than Equation 3, where only the expected safety cost is constrained:
$\min_{\pi}$ $\mathbb{E}_{\pi}\Big{[}\sum_{t}\gamma^{t}\mathcal{C}_{\text{task}}(\{\mathbf{x},\mathbf{y}_{<t}\},\text{y}_{t})\Big{]}$ (9)
[…]
While the formulation in Equation 9 is “stronger” than Equation 3, solving for the augmented MDP formulation with objective as Equation 7 can yield a policy satisfying the above almost sure constraints. We formally state this result in Theorem 4.1 and relegate the proof to Appendix B.3. {restatable}thmthmalmostsure (Almost Sure Safety) Consider an augmented MDP with cost function $\tilde{\mathcal{C}}_{\text{task}}^{\infty}$. Suppose an optimal policy exists $\pi^{\star}$ solving Equation 7 (see Theorem 4.1) with a finite cost, then $\pi^{\star}$ is an optimal policy for Equation 9, i.e., $\pi^{\star}$ is safe with probability approaching one or almost surely.
4.2 Algorithm and Practical Implementation
Building on our theoretical framework, we propose a search algorithm with two approaches: i) pre-training a small latent-space critic for cases where costs are available only for complete trajectories and ii) directly leveraging intermediate costs for search optimization.
Training a Latent-Space Critic. We make the usual assumption that trajectories terminate at a maximum length $T$. In this case, the value function simplifies to become: $\bar{V}^{n}(\mathbf{h}_{t},\mathbf{o}_{t},\text{z}_{t})=\mathbb{E}_{\bar{\pi}}[\gamma^{T}\bar{c}_{\text{task}}({\mathbf{h}}_{T},{\mathbf{o}}_{T})]$ if there is safety budget left, i.e., if $\text{z}_{T}>0$, or $n$ if $\text{z}_{T}\leq 0$, where $\bar{c}_{\text{task}}({\mathbf{h}}_{t},{\mathbf{o}}_{t})=c_{\text{task}}([\mathbf{x},\mathbf{y}_{\leq t}])$ in the latent MDP.
Hence, it is sufficient to predict: the sign of $\text{z}_{T}$ and the value of $\gamma^{T}\bar{c}_{\text{task}}({\mathbf{h}}_{T},{\mathbf{o}}_{T})$ to assess the quality of a state. We estimate those through Monte Carlo (MC) sampling. Specifically, we generate multiple trajectories from the initial state ($\mathbf{h}_{0},\mathbf{o}_{0},\text{z}_{0})$ using the reference policy, and compute the mean terminal cost, the sign of $z_{T}$ to serve as targets for the critic training. The usual alternative to MC sampling is Temporal Difference (TD) learning, where the critic is updated based on the difference between the current estimate and a bootstrapped estimate from the next state. However, MC sampling offers two advantages: i) it simplifies training by using separate supervised signals for quality and safety, unlike TD, which combines both, and ii) it allows dynamic adjustment of $n$ without retraining.
We train a critic network with two heads by sampling responses from the base model and scoring them using the cost function. We define $\mathcal{J}_{1}$ as the binary cross-entropy for predicting the sign of $\text{z}_{T}$ and $\mathcal{J}_{2}$ as the mean squared error for predicting $\gamma^{T}\bar{c}_{\text{task}}({\mathbf{h}}_{T},{\mathbf{o}}_{T})$. Our critic training minimizes: $\mathcal{J}(\mathbf{\theta})=\mathbb{E}_{\bar{\pi}}\Big{[}\sum_{t=1}^{T}\mathcal{J}_{1}\left(f^{1}_{\mathbf{\theta}}({\mathbf{h}}_{t},{\mathbf{o}}_{t},\text{z}_{t}),\text{z}_{T}>0\right)+\mathcal{J}_{2}\left(f^{2}_{\mathbf{\theta}}({\mathbf{h}}_{t},{\mathbf{o}}_{t},\text{z}_{t}),\gamma^{T}\bar{c}_{\text{task}}({\mathbf{h}}_{T},{\mathbf{o}}_{T})\right)\Big{]}$, where $f^{1}_{\mathbf{\theta}}$ and $f^{2}_{\mathbf{\theta}}$ are the heads of our parameterized critic.
Search method. We build on the beam search strategy from (Mudgal et al., 2023; Li, 2025) wherein we sequentially sample $N$ beams of $d$ tokens into a set $B$ from the pre-trained model and choose $K$ beams with the highest scores as possible continuations of the prompt (see Algorithm 1 in Appendix C). This step ensures that we focus on the most promising continuations. The goal of the scoring function is to balance the immediate task cost and the predicted future task cost while ensuring safety. This is repeated until we complete trajectories. Given a token trajectory $\text{y}_{t:t+d}$, we present a scoring function $\text{E}_{\text{critic}}$ where we assume that we cannot evaluate intermediate answers with the cost functions. However, when immediate safety costs are available, a simpler scoring function, see Appendix C. We define $\text{E}_{\text{critic}}$ as:
$\text{E}_{\text{critic}}(\text{y}_{t:t+d})=$$\begin{cases}\gamma^{T}\bar{c}_{\text{task}}(\cdot)&t+d=T\text{ and }\text{z}_{t+d}>0\\ n&t+d=T\text{ and }\text{z}_{t+d}\leq 0\\ f^{2}_{\mathbf{\theta}}(\cdot)&f^{1}_{\mathbf{\theta}}(\cdot)>0.5\\ n&\text{otherwise}.\end{cases}$
This $\text{E}_{\text{critic}}$ scoring function evaluates token sequences by balancing safety and task performance. At the final step ($t+d=T$), it assigns a score based on the task cost $\mathcal{C}_{\text{task}}$ if safety constraints are met ($\text{z}_{t+d}>0$); otherwise, it applies a high penalty $n$. For intermediate steps, it relies on a trained critic. If the critic confidently predicts safety ($f_{\mathbf{\theta}}^{1}(\mathbf{h}_{t+d},\mathbf{o}_{t+d},\text{z}_{t+d})>0.5$), it uses the estimated future cost ($f_{\mathbf{\theta}}^{2}(\mathbf{h}_{t+d},\mathbf{o}_{t+d},\text{z}_{t+d})$); otherwise, it assigns the penalty $n$ as a conservative safeguard.
Figure 2: Trade-offs between safety, reward, and inference time for BoN, ARGS, RECONTROL, Beam Search, and InferenceGuard on the PKU-SafeRLHF test dataset, evaluated for Alpaca-7B (top) and Beaver-v3-7B (bottom). Reward is the average score evaluated by the reward model, safety rate is the percentage of tasks completed within budget $d$, and inference time is the average per-task duration. The left column (reward vs. safety) and right column (inference time vs. safety) categorize methods by performance. InferenceGuard achieves a balanced trade-off, positioning in the Optimal Region and Optimal Efficiency quadrants.
Sampling Diversity. Finally, if the right selection strategy can guarantee that we will converge on a safe solution, it does not consider how many samples would be necessary. To increase the search speed, we introduce a diversity term in the sampling distribution when no safe samples were found based on the token frequency of failed beams. We denote $F$ as the frequency matrix counting the tokens we previously sampled from $t$ to $t+d$. Instead of sampling from $\text{SoftMax}(\mathbf{W}\mathbf{o}_{t})$, we sample from $\text{SoftMax}(\mathbf{W}\mathbf{o}_{t}-\text{n}_{2}({F_{t}>0}))$ where $\text{n}_{2}({F_{t}>0})$ is a vector where each component $j$ is $\text{n}_{2}$ if $F_{t,j}>0$ and 0 otherwise. The addition of $\text{n}_{2}({F_{t}>0})$ disables the possibility of sampling the same token at the same position observed in unsuccessful beams, thus increasing diversity.
It is worth noting that as we sample from the reference LLM and rank responses directly or via the critic, block sampling ensures a small Kullback-Leibler (KL) divergence from the original LLM without explicitly adding a KL regularizer into our objective, preserving coherence and natural flow; see (Mudgal et al., 2023).
5 Experiments
Figure 3: Reward and cost distributions on PKU-SafeRLHF test tasks using Alpaca-7B (top) and Beaver-v3 (bottom) as base models. The left y-axis shows reward distribution. The right y-axis shows the maximum cumulative cost. InferenceGuard outperforms others, achieving higher rewards with lower costs across both models.
Baselines. We evaluate the helpfulness (task cost) and harmlessness (safety cost) of our method on both non-safety-aligned and safety-aligned base models, Alpaca-7B Taori et al. (2023) and Beaver-7B (Ji et al., 2024b). We compare InferenceGuard to the following state-of-the-art test-time alignment methods: Best-of-N (BoN), beam search, ARGS (Khanov et al., 2024) and RECONTROL (Kong et al., 2024). These test-time alignment methods were originally designed to maximize rewards without considering safety. To ensure a fair and meaningful comparison, we extend them to incorporate safe alignment for LLMs. We incorporate a Lagrangian-based approach for all baselines that penalizes safety violations, ensuring a balanced trade-off between task performance and safety. This helps us evaluate our performance against other algorithms and highlights the importance of safety augmentation instead of a Lagrangian approach for effectively balancing rewards and constraint satisfaction.
For beam search and Best-of-N (BON), we adopt a Lagrangian approach to select solutions with $c_{\text{task}}+\lambda\mathcal{C}_{\text{safety}}$ where $\lambda$ is the Lagrangian multiplier. Similarly, we extend ARGS so that token selection follows: $-\omega\pi(t|\cdot)+c_{\text{task}}+\lambda\mathcal{C}_{\text{safety}}$, with $\omega$ adjusting the influence of the reference policy. We also considered state augmentation for ARGS and RE-control but found it ineffective. Since these methods decode token-by-token, they cannot recover once $\text{z}_{t}$ flips the sign, and before that, $\text{z}_{t}$ has no influence. Thus, we excluded it from our evaluation. To further strengthen our comparison, we introduce safety-augmented versions of BoN and beam search as additional baselines.
Datasets. We evaluate all the above methods on the PKU-SafeRLHF dataset (Ji et al., 2024b), a widely recognized benchmark for safety assessment in the LLM literature. This dataset contains 37,400 training samples and 3,400 testing samples. Of course, our training samples are only used to train the critic value network. Here, we construct a dataset by generating responses using the base models with five sampled trajectories for each PKU-SafeRLHF prompt from the training set.
Evaluation Metrics. We assess the performance using several metrics: the Average Reward is computed using the reward model of (Khanov et al., 2024) as $-c_{\text{task}}$ on the complete response to reflect helpfulness, where a higher reward indicates better helpfulness; the Average Cost is evaluated with the token-wise cost model from (Dai et al., 2023) as $\mathcal{C}_{\text{safety}}$, indicating harmfulness, with higher cost values reflecting more resource-intensive or harmful outputs; the Safety Rate is defined as the proportion of responses where the cumulative cost does not exceed the safety budget $z_{t=0}=10$, and is formally given by Safety $\text{Safety Rate}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}(\mathcal{C}^{i}_{\text{test}}\leq z_{t=0})$, where $N$ is the total number of responses; and Time refers to the inference time taken to generate a response in seconds.
Results. We present our main results in Figure 2 and additional ones in Table 2 in Appendix D. InferenceGuard achieves the highest safety rates with both models (91.04% and 100% for Alpaca and Beaver respectively). With Beaver, our method dominates the Patero front, achieving the highest rewards without any unsafe responses. Although Lagrangian methods can have a reasonable average cost, they all fail to satisfy the safety constraints. In the meantime, they are too safe on already safe answers, hindering their rewards. The RE-control intervention method underperforms for both models in our setting. ARGS can provide safe answers but with very poor rewards because most answers are very short to avoid breaking the safety constraint. Best-of-N and beam search with augmented safety came after us. However, Best-of-N cannot leverage intermediate signal, and beam search is blind to its previous mistakes, yielding more unsafe answers.
Figure 3 provides a better view of the reward and safety distributions. The figure shows that InferenceGuard consistently achieves higher rewards while maintaining low cumulative costs, outperforming other methods across both models. Its maximum cumulative cost stays just under the safety budget while maximizing the reward as suggested by our theoretical contributions. Finally, we observe that the trained critic helps to better guide the search on intermediate trajectories in both settings. We also compare generated answers in Appendix D qualitatively.
6 Related Work
Safe RL: Safe RL employs the cMDP framework, formulated in (Altman, 1999). Without prior knowledge, (Turchetta et al., 2016; Koller et al., 2018; Dalal et al., 2018) focus on safe exploration, while with prior knowledge, (Chow et al., 2018; 2019; Berkenkamp et al., 2017) learn safe policies using control techniques. (Achiam et al., 2017; Ray et al., 2019; Stooke et al., 2020) enforce safety constraints via Lagrangian or constrained optimization which can often lead to suboptimal safety-reward trade-offs. Instead, we extend safety state augmentation Sootla et al. (2022) to LLMs and latent MDPs to ensure almost sure inference time safety.
LLM alignment and safety: Pre-trained LLMs are often aligned to specific tasks using RL from Human Feedback (RLHF), where LLMs are fine-tuned with a learned reward model (Stiennon et al., 2020; Ziegler et al., 2019; Ouyang et al., 2022) or directly optimized from human preferences (Rafailov et al., 2023; Azar et al., 2023; Zhao et al., 2023; Tang et al., 2024; Song et al., 2024; Ethayarajh et al., 2024). (Bai et al., 2022; Ganguli et al., 2022) first applied fine-tuning in the context of safety, and (Dai et al., 2023) proposed safe fine-tuning via Lagrangian optimization. Other safety-focused methods such as (Gundavarapu et al., 2024; Gou et al., 2024; Hammoud et al., 2024; Hua et al., 2024; Zhang et al., 2024b; Guo et al., 2024; Xu et al., 2024; Wei et al., 2024; Li et al., 2025a) are either orthogonal, handle a different problem to ours, or can not ensure almost sure safety during inference.
Inference time alignment: Inference-time alignment of LLMs is often performed through guided decoding, which steers token generation based on rewards (Khanov et al., 2024; Shi et al., 2024; Huang et al., 2024) or a trained value function Han et al. (2024); Mudgal et al. (2023); Kong et al. (2024). Other safety-focused inference-time methods include (Zhong et al., 2024b; Banerjee et al., 2024; Niu et al., 2024; Wang et al., 2024a; Zeng et al., 2024; Zhao et al., 2024). Compared to those methods, we are the first to theoretically guarantee almost sure safe alignment with strong empirical results. Operating in the latent space enables us to train smaller, inference-efficient critics while optimally balancing rewards and safety constraints without introducing extra parameters, e.g., Lagrangian multipliers.
We detail other related works extensively in Appendix A.
7 Conclusion
We introduced InferenceGuard, a novel inference-time alignment method that ensures large LLMs generate safe responses almost surely—i.e., with a probability approaching one. We extended prior safety-augmented MDP theorems into the latent space of LLMs and conducted a new analysis. Our results demonstrated that InferenceGuard significantly outperforms existing test-time alignment methods, achieving state-of-the-art safety versus reward tradeoff results. In the future, we plan to improve the algorithm’s efficiency further and generalize our setting to cover jailbreaking. While our method is, in principle, extendable to jailbreaking settings, we aim to analyze whether our theoretical guarantees still hold.
Broader Impact Statement
This work contributes to the safe and responsible deployment of large language models (LLMs) by developing InferenceGuard, an inference-time alignment method that ensures almost surely safe responses. Given the increasing reliance on LLMs across various domains, including healthcare, education, legal systems, and autonomous decision-making, guaranteeing safe and aligned outputs is crucial for mitigating misinformation, bias, and harmful content risks.
To further illustrate the effectiveness of our approach, we have included additional examples in the appendix demonstrating that our method successfully produces safe responses. These examples were generated using standard prompting with available large language models LLMs. Additionally, we have added a warning at the beginning of the manuscript to inform readers about the nature of these examples. Our primary motivation for this addition is to highlight the safety improvements achieved by our method compared to existing alternatives. We do not foresee these examples being misused in any unethical manner, as they solely showcase our model’s advantages in ensuring safer AI interactions. Finally, we emphasize that our method is designed specifically to enhance AI safety, and as such, we do not anticipate any potential for unethical applications.
InferenceGuard enhances the scalability and adaptability of safe AI systems by introducing a formally grounded safety mechanism that does not require model retraining while reducing the resource costs associated with traditional RLHF methods. The proposed framework advances AI safety research by providing provable safety guarantees at inference time, an area that has received limited attention in prior work.
While this method significantly improves safety in LLM outputs, it does not eliminate all potential risks, such as adversarial manipulation or emergent biases in model responses. Future work should explore robustness to adversarial attacks, contextual fairness, and ethical considerations in deploying safety-aligned LLMs across different cultural and regulatory landscapes. Additionally, transparency and accountability in AI safety mechanisms remain essential for gaining public trust and ensuring alignment with societal values. This work aims to empower developers and policymakers with tools for ensuring safer AI deployment while contributing to the broader conversation on AI ethics and governance.
Acknowledgments
The authors would like to thank Rasul Tutunov, Abbas Shimary, Filip Vlcek, Victor Prokhorov, Alexandre Maraval, Aivar Sootla, Yaodong Yang, Antonio Filieri, Juliusz Ziomek, and Zafeirios Fountas for their help in improving the manuscript. We also would like to thank the original team from Pekin University, who made their code available and ensured we could reproduce their results.
Appendix A Additional Related Work
Safe RL: Safe RL employs the cMDP framework (Altman, 1999) to enforce safety constraints during exploration and policy optimization. When no prior knowledge is available, methods focus on safe exploration (Turchetta et al., 2016; Koller et al., 2018; Dalal et al., 2018; Wachi & Sui, 2018; Bharadhwaj et al., 2020). With prior knowledge, such as environmental data or an initial safe policy, methods learn safe policies using control techniques like Lyapunov stability (Chow et al., 2018; 2019; Berkenkamp et al., 2017; Ohnishi et al., 2019) and reachability analysis (Cheng et al., 2019; Akametalu et al., 2014; Dean et al., 2019; Fisac et al., 2019). Safety constraints are enforced via Lagrangian or constrained optimization methods (Achiam et al., 2017; Ray et al., 2019; Stooke et al., 2020; Yang et al., 2019; Ding et al., 2020; Ji et al., 2024b), but can often lead to suboptimal safety-reward trade-offs. In contrast, our approach extends safety state augmentation Sootla et al. (2022) to LLMs and latent MDPs to ensure almost sure inference time safety without relying on Lagrangian multipliers.
LLM alignment and safety: Methods for aligning pre-trained LLMs with task-specific data include prompting, guided decoding, and fine-tuning. Among fine-tuning methods, RL from Human Feedback (RLHF) has proven effective, where LLMs are fine-tuned with a learned reward model (Stiennon et al., 2020; Ziegler et al., 2019; Ouyang et al., 2022) or directly optimized from human preferences (Rafailov et al., 2023; Azar et al., 2023; Zhao et al., 2023; Tang et al., 2024; Song et al., 2024; Ethayarajh et al., 2024; Ramesh et al., 2024). Recent works have explored fine-tuning for helpful and harmless responses (Bai et al., 2022; Ganguli et al., 2022), while (Dai et al., 2023) introduced a safe RL approach incorporating safety cost functions via Lagrangian optimization, requiring model weight fine-tuning. Other safety-focused methods, including machine unlearning (Gundavarapu et al., 2024), safety pre-aligned multi-modal LLMs (Gou et al., 2024), safety-aware model merging (Hammoud et al., 2024), prompting-based safety methodologies (Hua et al., 2024), test-time controllable safety alignment (Zhang et al., 2024b), defenses against adversarial attacks and jailbreaking (Guo et al., 2024; Qi et al., 2024; Xu et al., 2024), identifying safety criticial regions in LLMs (Wei et al., 2024), safety preserved LoRA fine-tuning (Li et al., 2025a), alignment using correctional residuals between preferred and dispreferred answers using a small model (Ji et al., 2024a), and identifying safety directions in embedding space (Feng et al., 2024).Those methods are either orthogonal, handle a different problem to ours, or can not ensure almost sure safety during inference.
Inference time alignment: The closest literature to ours is inference-time alignment. Those methods offer flexible alternatives to fine-tuning LLMs, as they avoid modifying the model weights. A common approach is guided decoding, which steers token generation based on a reward model. In particular, (Khanov et al., 2024; Shi et al., 2024; Huang et al., 2024) perform this guided decoding through scores from the reward model whereas Han et al. (2024); Mudgal et al. (2023); Kong et al. (2024) use a value function that is trained on the given reward model. These inference-time alignment methods build on previous works like (Yang & Klein, 2021; Arora et al., 2022; Krause et al., 2021; Kim et al., 2023; Meng et al., 2022; Peng et al., 2019), which guide or constrain LLMs towards specific objectives. Other safety-focused inference-time methods include, reverse prompt contrastive decoding (Zhong et al., 2024b), adjusting model hidden states combined with guided decoding (Banerjee et al., 2024), soft prompt-tuned detoxifier based decoding (Niu et al., 2024), jail-break value decoding (Wang et al., 2024a), speculating decoding using safety classifier (Zeng et al., 2024), and opposite prompt-based contrastive decoding (Zhao et al., 2024). Compared to those methods, we are the first to achieve almost sure safe alignment with strong empirical results. Operating in the latent space enables us to train smaller, inference-efficient critics while optimally balancing rewards and safety constraints (see Section 5) without introducing additional parameters like Lagrangian multipliers.
Appendix B Theoretical Analysis
For our theoretical results, we consider a similar setup to that of Sootla et al. (2022); Hernández-Lerma & Muñoz de Ozak (1992) but with a discrete action space. Consider an MDP $M=\{S,A,P,c,\gamma_{c}\}$ with a discrete, non-empty, and finite action set for each $s$ defined as $A(s)$. The set
$K=\{(s,a)\mid s\in S,a\in A(s)\}$
defines the admissible state-action pairs and is assumed to be a Borel subset of $S\times A$. A function $u$ is inf-compact on $K$ if the set $\{a\in A(s)\mid u(s,a)\leq r\}$ is compact for every $s\in S$ and $r\in\mathbb{R}$. Note that, since the action space is finite and discrete every function $u$ is inf-compact on $K$. A function $u$ is lower semi-continuous (l.s.c.) in $S$ if for every $s_{0}\in S$ we have
$\liminf_{s\to s_{0}}u(s)\geq u(s_{0}).$
Let $L(S)$ denote the class of all functions on $S$ that are l.s.c. and bounded from below.
For a given action, $a\in A(s)$, a distribution $P(y\mid s,a)$ is called weakly continuous w.r.t. $s$, if for any function $u(s)$, continuous and bounded w.r.t. $s$ on $S$, the map
$(s,a)\mapsto\int_{S}u(y)P(dy\mid s,a)$
is continuous on $K$ for a given $a$.
We also make the following assumptions:
B1. The function $c(s,a)$ is bounded, measurable on $K$, nonnegative, lower semi-continuous w.r.t. $s$ for a given $a\in A(s)$;
B2. The transition law $P$ is weakly continuous w.r.t. $s$ for a given $a\in A(s)$;
B3. The set-value function map $s:A(s)$ satisfies the following, $\forall s_{0}\in\mathcal{S}$, there exists a $\epsilon>0$, such that $\forall x$ satisfying $\|x-x_{0}\|\leq\epsilon$, $A(x)=A(x_{0})$
Note that, the assumptions B1-B3, share a similar essence to that of the Assumptions 2.1-2.3 in Hernández-Lerma & Muñoz de Ozak (1992) and B1-B3 in Sootla et al. (2022) but suited for a discrete action space. In particular, Assumption B3, is similar to the lower semi continuity assumption on the set-value function map $A(s)$ taken in Sootla et al. (2022); Hernández-Lerma & Muñoz de Ozak (1992) but modified for a discrete action space.
Our first goal is to recreate Hernández-Lerma & Muñoz de Ozak (1992, Lemma 2.7) for our discrete action setting. Let $\Pi$ denote the set of functions from $S\to A$.
Lemma 1.
  • (a)
    If Assumption B3 holds and $v(s,a)$ is l.s.c. w.r.t. $s$ for any given $a\in A(s)$ and bounded from below on $K$, then the function
    $v^{*}(s):=\inf_{a\in A(s)}v(s,a)$
    belongs to $L(S)$ and, furthermore, there is a function $\pi\in\Pi$ such that
    $v^{*}(s)=v(s,\pi(s))\quad\forall s\in S.$
  • (b)
    If the Assumptions B1-B3 hold, and $u\in L(S)$ is nonnegative, then the (nonnegative) function
    $u^{*}(s):=\inf_{a\in A}\left[c(s,a)+\int_{S}u(y)P(dy\mid s,a)\right]$
    belongs to $L(S)$, and there exists $\pi\in\Pi$ such that
    $u^{*}(s)=c(s,\pi(s))+\int_{S}u(y)P(dy\mid s,\pi(s))\quad\forall s\in S.$
  • (c)
    For each $n=0,1,\dots$, let $v_{n}$ be a l.s.c. function, bounded from below. If $v_{n}\to v_{0}$ as $n\to\infty$, then
    $\lim_{n\to\infty}\inf_{a\in A(s)}v_{n}(s,a)=\inf_{a\in A(s)}v_{0}(s,a)\quad\forall s\in S.$
Proof.
For part a)
We have $v(s,a)$ is l.s.c. w.r.t. $s$ for any given $a$. This implies from the definition of lower semi -continuity for any $s_{0}\in S$ and $a\in A(s_{0})$, if $v(s_{0},a)>y$, then there exists a $\epsilon>0$, s.t. $\forall s$ satisfying $\|s-s_{0}\|\leq\epsilon$, $v(s,a)>y$.
Assume for some $s_{0}\in S$ and $y$, the function $\inf_{a\in A(s_{0})}v(s_{0},a)$ satisfies,
$\inf_{a\in A(s_{0})}v(s_{0},a)>y\quad\Rightarrow v(s_{0},a)>y\quad\forall a\in A(s_{0})$ (10)
Using Assumption B3, we have $A(s)=A(s_{0})$ for $\|s-s_{0}\|\leq\epsilon$. Moreover, using the fact that $v(s,a)$ is l.s.c. at a given $a$, we have if $v(s_{0},a)>y$, $\forall a\in A(s_{0})$ we have,
$v(s,a)>y\quad\forall\|s-s_{0}\|\leq\epsilon,\forall a\in A(s)$ (11)
Since, this holds for all $a\in A(s)$, it also holds for
$\inf_{a\in A(s)}v(s,a)>y\quad\forall\|s-s_{0}\|\leq\epsilon$ (12)
This proves the lower semi continuity of $v^{*}(s)=\inf_{a\in A(s)}v(s,a)$. Further, due to the discrete nature of $A(s)$, the $\inf_{a\in A(s)}v(s,a)$ is always attained by an action $\pi(s)\in A(s)$. Hence, there exists a function $\Pi:S->A(s)$ s.t.
$v^{*}(s)=\inf_{a\in A(s)}v(s,a)=v(s,\pi(s))\quad\forall s\in S.$ (13)
For part b), note that $c(s,a)+\int u(y)P(dy\mid s,a)$ is l.s.c. w.r.t. $s$ for an given $a$, based on Assumptions B1-B2. Hence, using part a) we have,
$u^{*}(s):=\inf_{a\in A}\left[c(s,a)+\int_{S}u(y)P(dy\mid s,a)\right]=c(s,\pi(s))+\int_{S}u(y)P(dy\mid s,\pi(s))\in L(S)\quad\forall s\in S$ (14)
for some $f\in\Pi$.
For part c), we begin by defining $l(s)=lim_{n\to\infty}\inf_{a\in A(s)}v_{n}(s,a)$. Note that, since $\{v_{n}\}$ is an increasing sequence, we have for any $n$
$\inf_{a\in A(s)}v_{n}(s,a)\leq\inf_{a\in A(s)}v_{0}(s,a)$ (15)
This implies,
$l(s)\leq\inf_{a\in A(s)}v_{0}(s,a)=v_{0}^{*}(s)$ (16)
Next, we define for any $s\in S$,
$A_{n}:=\{a\in A(s)|v_{n}(s,a)\leq v_{0}^{*}(s)\}$ (17)
We note that $A_{n}$ are compact sets as $A$ is finite and discrete. Further, note that $A_{n}$ is a decreasing sequence converging to $A_{0}$ (compact, decreasing and bounded from below by $A_{0}$). Also, note that
$A_{1}\supset A_{2}\supset A_{3}\cdots\supset A_{0}$ (18)
We consider the sequence $\{a_{n}\}$ where $a_{n}\in A_{n}$ and $a_{n}$ satisfies,
$v_{n}(s,a_{n})=\inf_{a\in A(s)}v_{n}(s,a)\leq\inf_{a\in A(s)}v_{0}(s,a)\leq v_{0}^{*}(s)$ (19)
This sequence $\{a_{n}\}$ belongs to the compact space $\cup_{n=1}^{\infty}A_{n}=A_{1}$, hence it has convergent subsequence $\{a_{n_{i}}\}$ converging to $\cup_{n=1}^{\infty}A_{n}=A_{1}$.
$a_{n_{i}}$ $\in A_{n_{i}}=\cap_{n\leq n_{i}}A_{n}$ (20)
$a_{0}$ $\in\cap_{n\leq\infty}A_{n}=A_{0}$ (21)
Since, the converging sequence $a_{n_{i}}\to a_{0}$ belongs to the discrete, compact space, there exisits a $N_{i}$, such that for all $n_{i}\geq N_{i}$, $a_{n_{i}}=a_{0}$. Further, using the increasing nature of $v_{n}$, we have,
$v_{n_{i}}(s,a_{n_{i}})\geq v_{n}(s,a_{n_{i}})\quad\forall n_{i}\geq n$ (22)
As $i\to\infty$, this implies,
$\lim_{i\to\infty}v_{n_{i}}(s,a_{n_{i}})\geq v_{n}(s,a_{0})$ (23)
$\lim_{i\to\infty}\inf_{a\in A(s)}v_{n_{i}}(s,a)\geq v_{n}(s,a_{0})$ (24)
$l(s)\geq v_{n}(s,a_{0})$ (25)
As $v_{n}\to v_{0}$, $l(s)\geq v_{0}(s,a_{0})=v_{0}^{*}(s)$.
B.1 Optimality Equation
Next, we characterize the solution to the Bellman (Optimality) equation. We begin by recalling the Bellman operator:
$Tv(s)=\min_{a\in A(s)}\left\{c(s,a)+\gamma\int v(y)P(dy\mid s,a)\right\}.$
To state our next result we introduce some notation: Let $L(S)^{+}$ be the class of nonnegative and l.s.c. functions on $S$, and for each $u\in L(S)^{+}$ by Lemma 1(b), the operator $T$ maps $L(S)^{+}$ into itself. We also consider the sequence $\{v_{n}\}$ of value iteration (VI) functions defined recursively by
$v_{0}(S):=0,\quad\text{and}\quad v_{h}:=Tv_{h-1}\quad\text{for}\quad h=1,2,\dots$
That is, for $h\geq 1$ and $s\in S$,
$v_{h}(s):=\min_{a\in\mathcal{A}(s)}\left(c(s,a)+a\int_{S}v_{h-1}(y)P(dy\mid s,a)\right).$
Note that, by induction and Lemma 1(b) again, $v_{h}\in L(S)^{+}$ for all $h\geq 0$. From elementary Dynamic Programming Bertsekas (1987); Bertsekas & Shreve (1996); Dynkin & Yushkevich (1979), $v_{h}(s)$ is the optimal cost function for an $h$-stage problem (with ”terminal cost” $v_{0}(s)=0$) given $s_{0}=s$; i.e.,
$v_{h}(s)=\inf_{\pi}V_{h}(\pi,s),$
where, $\Pi$ is the set of policies and $V_{H}(\pi,s)$ denotes the value function for the $H-$stage problem:
$V_{H}(\pi,s_{0})=\mathbb{E}_{\pi}\left[\sum_{h=0}^{H-1}\gamma^{h}c(s_{h},a_{h})\right].$
Here, $\mathbb{E}_{\pi}$ stands for the expectation with actions sampled according to the policy $\pi$ and the transitions $P$. For $H\to\infty$, let the value functions be denoted as follows:
$V(\pi,s_{0})=\mathbb{E}_{\pi}\left[\sum_{h=0}^{\infty}\gamma^{h}c(s_{h},a_{h})\right],$
and
$V^{*}(s)=\inf_{\pi}V(\pi,s).$
We want to prove similar results to that of Hernández-Lerma & Muñoz de Ozak (1992, Theorem 4.2) on the optimality of the Bellman operator, however in the discrete action setting. In particular, we want to show the following theorem
Theorem 2.
Suppose that Assumptions B1-B3 hold, then:
  • (a)
    $v_{h}\to V^{*}$; hence
  • (b)
    $V^{*}$ is the minimal pointwise function in $L(S)^{+}$ that satisfies
    $V^{*}=TV^{*}$
Proof.
We follow a similar proof strategy to that of Hernández-Lerma & Muñoz de Ozak (1992, Theorem 4.2).
To begin, note that the operator $T$ is monotone on $L(S)^{+}$, i.e., $u>v$ implies $Tu>Tv$. Hence $\{v_{h}\}$ forms a nondecreasing sequence in $L(S)^{+}$ and, therefore, there exists a function $u\in L(S)^{+}$ such that $v_{h}\to u$. This implies (by the Monotone Convergence Theorem) that
$c(s,a)+a\int v_{h-1}(y)P(dy\mid s,a)\to c(s,a)+a\int u(y)P(dy\mid s,a),$
Using Lemma 1(c), and $v_{h}=\inf_{a\in A(s)}\{c(s,a)+a\int v_{h-1}(y)P(dy\mid s,a)\}$ yields
$\lim_{h\to\infty}\inf_{a\in A(s)}\{c(s,a)+a\int v_{h-1}(y)P(dy\mid s,a)\}$ $=\inf_{a\in A(s)}\{c(s,a)+a\int u(y)P(dy\mid s,a)\},$
$\lim_{h\to\infty}v_{h}$ $=Tu,$
$u$ $=Tu.$
This shows $v_{h}\to u$, such that $u\in L(S)^{+}$ satisfies the Optimality equation.
Next, we want to show $u=V^{*}$. Using that $u\geq Tu$, and by Lemma 1(b), we have that there exists $\pi\in\Pi$, a stationary policy that satisfies
$u(s)$ $\geq\inf_{a\in A(s)}\{c(s,a)+a\int u(y)P(dy\mid s,a)\}\quad\forall s$ (26)
$\geq c(x,\pi)+\alpha\int u(y)P(dy\mid s,\pi)\quad\forall s.$ (27)
Applying the $T$ operator iteratively, we have
$u(s)$ $\geq T^{H}u(s)\quad\forall s,H$ (28)
$\geq\mathbb{E}_{\{s_{h}\},\pi}[\sum_{h=0}^{H-1}\alpha^{h}c(s_{h},\pi)]+\alpha^{H}\int u(y)P^{H}(dy\mid s,\pi)\quad\forall s,H,$ (29)
where $P^{H}(B\mid s,\pi)=P(\{s_{H}\in B\})$ denotes the $H$-step transition probability of the Markov chain $\{s_{h}\}$ (see Hernández-Lerma & Muñoz de Ozak (1992, Remarks 3.1,3.2)). Therefore, since $u$ is nonnegative,
$u(s)$ $\geq\mathbb{E}_{\{s_{h}\},\pi}[\sum_{h=0}^{H-1}\alpha^{h}c(s_{h},\pi)]\quad\forall s,H,$ (30)
Letting $H\to\infty$, we obtain
$u(s)\geq V(\pi,s)\geq V^{*}(s)\quad\forall s.$
Next, note that,
$v_{h}(s)$ $=\inf_{\pi\in\Pi}V_{h}(\pi,s)$ (31)
$\leq V_{h}(\pi,s)\quad\forall s,h,\pi$ (32)
and letting $h\to\infty$, we get
$u(s)\leq V(\pi,s)\quad\forall s,\pi.$
This implies $u(s)\leq V^{*}(s)$. We have thus shown that $u=V^{*}$.
Further, if there is another solution $u^{\prime}$ satisfying $u^{\prime}=Tu^{\prime}$, it holds that $u^{\prime}\geq V^{*}$, Hence, $V^{*}$ is the minimal solution.
B.2 Limit of a sequence of MDPs
Consider now a sequence of Markov Decision Processes (MDPs) $M_{n}=\{S,A,P,c_{n},\gamma_{n}\}$, where, without loss of generality, we write $c_{n},c_{\infty}$ and $M_{n},M_{\infty}$. Consider now a sequence of value functions $\{V_{n}^{*}\}_{n=0}^{\infty}$:
$V_{n}(\pi,s_{0})=\mathbb{E}^{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}c_{n}(s_{t},a_{t})\right],$
$V_{n}^{*}(s)=\inf_{\pi}V_{n}(\pi,s).$
The ”limit” value functions (with $n=\infty$) are still denoted as follows:
$V(\pi,s_{0})=\mathbb{E}^{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}c(s_{t},a_{t})\right],$
$V^{*}(s)=\inf_{\pi}V(\pi,s).$
We also define the sequence of Bellman operators
$T_{n}v(s)=\min_{a\in A(s)}\left\{c_{n}(s,a)+\gamma\int v(y)P(dy\mid s,a)\right\},$
$Tv(s)=\min_{a\in A(s)}\left\{c(s,a)+\gamma\int v(y)P(dy\mid s,a)\right\}.$
In addition to the previous assumptions, we make an additional one, while modifying Assumption B1:
  • B1’
    For each $n$, the functions $c_{n}(s,a)$ are bounded, measurable on $\mathcal{K}$, nonnegative, lower semi-continuous;
  • B4
    The sequence $\{c_{n}(s,a)\}_{n=0}^{\infty}$ is such that $c_{n}\uparrow c$.
For each $n$, the optimal cost function $V^{*}(s)$ is the bounded function in $L(S)^{+}$ that satisfies the Optimality equation in 2 :
$V_{n}^{*}=T_{n}V_{n}^{*},$
Theorem 3.
The sequence $V_{n}^{*}$ is monotone increasing and converges to $V^{*}$.
Proof.
We follow a similar proof strategy to that of Hernández-Lerma & Muñoz de Ozak (1992, Theorem 5.1)
To begin with, note that since $c_{n}\uparrow c$, it is clear that $V_{n}^{*}$ is an increasing sequence in $L(S)^{+}$, and therefore, there exists a function $u\in L(S)^{+}$ such that $V_{n}^{*}\to u$.
Moreover, from Lemma 1(c), letting $n\to\infty$, we see that $u=Tu$, i.e., $u$ satisfies the optimality equation. This implies that $u\geq V^{*}$, since, by Theorem 4, $V^{*}$ is the minimal solution in $L(X)^{+}$ to the optimality equation.
On the other hand, it is clear that $V_{n}^{*}\leq V^{*}$ for all $n$, so that $u\leq V^{*}$. Thus $u=V^{*}$, i.e., $U^{*}=V^{*}$.
B.3 Latent MDP Analysis
[…]
Proof.
We begin by comparing our assumptions to that of the assumptions B1-B4, closely aligned to those used in Hernández-Lerma & Muñoz de Ozak (1992); Sootla et al. (2022).
To prove a),b) of Definition 4.1 we need to verify that the latent MDP satisfying Assumptions A1-A2 also satisfies Assumptions B1’, B2-B4. According to Assumption A1, we consider bounded costs $\bar{\mathcal{C}}^{n}_{\text{task}}$ continuous w.r.t. state $(\mathbf{h},\mathbf{o},\text{z})$ for a given $y$ with discrete and finite action space ${\mathcal{V}}$, hence Assumptions B1’, B3, and B4 are satisfied. Assumptions B2 and A2 are identical. This proves a) and b).
For c), note that the state value function $V(\cdot)$ and latent space value function $\bar{V}(\cdot)$ w.r.t. policy $\bar{\pi}:\bar{{\mathcal{S}}}\rightarrow{\mathcal{V}}$ that acts on the latent space directly and on the original space as $\bar{\pi}(\phi(\cdot)):{\mathcal{S}}\rightarrow{\mathcal{V}}$ are related as follows:
$V(\bar{\pi}(\phi(\cdot)),\mathbf{s}_{0},\text{z}_{0})$ (33)
$=\mathbb{E}_{\begin{subarray}{c}\mathbf{s}_{t+1},\text{z}_{t+1}\sim{\mathcal{P}}(\cdot|\mathbf{s}_{t},\text{z}_{t},\text{y}_{t})\\ \text{y}_{t}\sim\bar{\pi}(\cdot|\phi(\mathbf{s}_{t}),\text{z}_{t})\end{subarray}}\Big{[}\sum_{t=0}^{\infty}\gamma^{t}\tilde{\mathcal{C}}^{n}_{\text{task}}(\mathbf{s}_{t},\text{z}_{t},\text{y}_{t})\Big{]}$ (34)
$=\mathbb{E}_{\begin{subarray}{c}\mathbf{s}_{t+1},\text{z}_{t+1}\sim{\mathcal{P}}(\cdot|\mathbf{s}_{t},\text{z}_{t},\text{y}_{t})\\ \text{y}_{t}\sim\bar{\pi}(\cdot|\phi(\mathbf{s}_{t}),\text{z}_{t})\end{subarray}}\Big{[}\sum_{t=0}^{\infty}\gamma^{t}\bar{\mathcal{C}}^{n}_{\text{task}}(\phi(\mathbf{s}_{t}),\text{z}_{t},\text{y}_{t})\Big{]}$ (35)
$=\mathbb{E}_{\begin{subarray}{c}\phi(\mathbf{s}_{t+1}),\text{z}_{t+1}\sim\bar{{\mathcal{P}}}(\cdot|\phi(\mathbf{s}_{t}),\text{z}_{t})\big{)}\\ \text{y}_{t}\sim\bar{\pi}(\cdot|\phi(\mathbf{s}_{t}),\mathbf{o}_{t})\end{subarray}}\Big{[}\sum_{t=0}^{\infty}\gamma^{t}\bar{\mathcal{C}}^{n}_{\text{task}}(\phi(\mathbf{s}_{t}),\text{z}_{t},\text{y}_{t})\Big{]}\Big{|}$ (36)
$=\mathbb{E}_{\begin{subarray}{c}\mathbf{h}_{t+1},\mathbf{o}_{t+1},\text{z}_{t+1}\sim\bar{{\mathcal{P}}}(\cdot|\mathbf{h}_{t},\mathbf{o}_{t},\text{z}_{t})\big{)}\\ \mathbf{h}_{t+1},\mathbf{o}_{t+1}=\phi(\mathbf{s}_{t+1})\\ \text{y}_{t}\sim\bar{\pi}(\cdot|\mathbf{h}_{t},\mathbf{o}_{t})\end{subarray}}\Big{[}\sum_{t=0}^{\infty}\gamma^{t}\bar{\mathcal{C}}^{n}_{\text{task}}(\mathbf{h}_{t},\mathbf{o}_{t},\text{z}_{t},\text{y}_{t})\Big{]}\Big{|}_{\mathbf{h}_{0},\mathbf{o}_{0}=\phi(s_{0})}$ (37)
$=\bar{V}(\bar{\pi},\mathbf{h}_{0},\mathbf{o}_{0},z_{0})\Big{|}_{\mathbf{h}_{0},\mathbf{o}_{0}=\phi(s_{0})}$ (38)
Hence, we can show $\bar{\pi}_{n}^{*}$ is optimal for $V_{n}(\cdot)$ as follows:
$V_{n}(\bar{\pi}_{n}^{*},\mathbf{s},\text{z})$ $=\bar{V}_{n}(\bar{\pi}_{n}^{*},\mathbf{h},\mathbf{o},\text{z})\Big{|}_{\mathbf{h},\mathbf{o}=\phi(s)}$ (39)
$=\min_{\bar{\pi}}\bar{V}_{n}(\bar{\pi},\mathbf{h},\mathbf{o},\text{z})\Big{|}_{\mathbf{h},\mathbf{o}=\phi(s)}$ (40)
$=\min_{\bar{\pi}}V_{n}(\bar{\pi}(\phi(\cdot)),\mathbf{s},\text{z})$ (41)
$=\min_{\pi}V_{n}(\pi,\mathbf{s},\text{z})$ (42)
Here, the minimization of $\pi$ is over set of all policies covered by $\bar{\pi}(\phi(\cdot))$ and we show that $\bar{\pi}_{n}^{*}(\phi(\cdot))$ is the optimal policy for the original space over this set of policies.
[…]
Proof.
We first note that if any trajectory with infinite cost has a finite probability, the cost would be infinite. Hence, all the trajectories with finite/positive probability have finite costs. This implies, the finite cost attained by $\pi^{\star}$ w.r.t. Equation 7 implies the satisfaction of constraints (Equation 9) almost surely (i.e. with probability 1). Combined with the fact that the policy $\pi^{\star}$ was obtained by minimizing the exact task cost as in Equation 9, Definition 4.1 follows. ∎
Appendix C Algorithmic details
Input: Initial prompt $s_{0}$, beam depth $d$, $N$ number of beams, max depth $D$, top $K$ beams, $M$ max retry
Initialize the beam $B\leftarrow\{s_{0}\}$
for $\frac{D}{K}$ iterations do
      $\forall i\in\{1,\dots d\},j\in\{1,...,|\mathcal{V}|\},F_{i,j}\leftarrow$ 0
       for $M$ rounds do
            $B_{new}\leftarrow$ Generate $N$ continuations of length $d$ from $B$ with $\pi_{\text{pen}}$ and $F$
[…]
             if $\exists i,E_{i}>0$ or last round then
                  $B\leftarrow B_{\text{new}}$ ;
                   break
             end if
            $F\leftarrow F+token\_frequency(B_{\text{new}})$
       end for
      Keep top $K$ beams in $B$ according to their scores E
end for
return Best trajectory in $B$ according to E
Algorithm 1 InferenceGuard
If the cost functions allow intermediate evaluations, we evaluate a beam $y_{t},\cdots,y_{t+d}$ but using our augmented cost function:
$\text{E}_{\text{inter}}(y_{t},\cdots,y_{t+d})=\begin{cases}\gamma^{T}\bar{c}_{\text{task}}({\bm{h}}_{t+d},{\bm{o}}_{t+d})&\text{z}_{t+d}>0\\ n&\text{otherwise}.\end{cases}$
When we only have a critic, we use:
$\text{E}_{\text{critic}}(y_{t},\cdots,y_{t+d})=$$\begin{cases}\gamma^{T}\bar{c}_{\text{task}}({\bm{h}}_{t+d})&t+d=T\text{ and }\text{z}_{t+d}>0\\ n&t+d=T\text{ and }\text{z}_{t+d}\leq 0\\ f^{2}_{\bm{\theta}}({\bm{h}}_{t+d},{\bm{o}}_{t+d},\text{z}_{t+d})&f^{1}_{\bm{\theta}}({\bm{h}}_{t+d},{\bm{o}}_{t+d},\text{z}_{t+d})>0.5\\ n&\text{otherwise}.\end{cases}$
If we can do both, as $\text{z}_{t}$ only decreases overtime, the critic head predicting the safety would only act as an additional filter. We introduce another hyper-parameter $\eta$ to balance the confidence in the second head of the critic predicting the future cost:
$\text{E}_{\text{mix}}(y_{t},\cdots,y_{t+d})=\begin{cases}\gamma^{T}\bar{c}_{\text{task}}(\mathbf{h}_{t+d},\bm{o}_{t+d})&t+d=T\text{ and }\text{z}_{t+d}>0\\ n&t+d=T\text{ and }\text{z}_{t+d}\leq 0\\ \gamma^{T}\bar{c}_{\text{task}}({\bm{h}}_{t+d},\bm{o}_{t+d})+\eta f^{2}_{\bm{\theta}}({\bm{h}}_{t+d},{\bm{o}}_{t+d},\text{z}_{t+d})&f^{1}_{\bm{\theta}}({\bm{h}}_{t+d},{\bm{o}}_{t+d},\text{z}_{t+d})>0.5\text{ and }\text{z}_{t+d}>0\\ n&\text{otherwise}.\end{cases}$
Appendix D Experimental Details
D.1 Experiment Setup
We conducted all experiments on NVIDIA V100 GPUs, utilizing Python 3.11.11 and CUDA 12.4, for model development and training. The experimental setup was based on the PKU-SaferRLHF dataset.Our experiments involved two primary models: the unsafe-aligned reproduced Alpaca-7B model [^6] https://huggingface.co/PKU-Alignment/alpaca-7b-reproduced (Taori et al., 2023) and the safe-aligned Beaver-v3-7B model (Ji et al., 2024b). We employed the reward model Khanov et al. (2024) [^7] https://huggingface.co/argsearch/llama-7b-rm-float32 and a cost model Ji et al. (2024b) [^8] https://github.com/PKU-Alignment/safe-rlhf during the training stage for critic network, test-time inference and evaluation stages.
D.2 Baselines and Hyper-parameter Settings
The baseline methods we compare against include BoN, Beam Search, RECONTROL (Kong et al., 2024), and ARGS (Khanov et al., 2024). For BoN, we use a strategy where the top N outputs are sampled, and the best result is selected based on a predefined criterion. Similarly, Beam Search explores multiple beams during the search process and selects the best output based on a beam-width parameter. In RECONTROL, an MLP network is employed as the value network to intervene in the decision-making process, guiding the generation through reinforcement learning (Kong et al., 2024). ARGS, on the other hand, implements a logits-based greedy token-wise search strategy, where tokens are generated sequentially based on the maximum likelihood of the next token (Khanov et al., 2024).
Given the limited research on safe alignment with inference-time methods, we adapt these baseline methods to enable safe inference, ensuring a fair comparison. To this end, we incorporate a Lagrangian multiplier term to control the balance between reward and cost, allowing for safe inference based on their open-source implementations. Notably, BoN and Beam Search utilize a form of blocking sampling, while ARGS and RECONTROL employ token sampling methods.
In our setup, we modify the inference process of the baseline methods with a Lagrangian multiplier by using the following score for token selection: $c_{\text{task}}+\lambda\mathcal{C}_{\text{safety}}$ where $\lambda$ is the Lagrangian multiplier, that controls the influence of the safety cost score $\mathcal{C}_{\text{safety}}$. We unify $\lambda=5$ and also the sampling parameters for baseline methods and InferenceGuard to ensure a fair comparison. Each method has its own specific settings, as follows:
Method Sample Width $(d)$ Num Samples $(N)$ Other Parameters
ARGS 1 20 N/A
RECONTROL 1 N/A n steps = 30, step size = 0.5
BoN 32 100, 200, 500 N/A
Beam Search 32 128, 256 D=128, K=$N/4$
InferenceGuard 32 128, 512 M=2, D=128, K=$N/4$
Table 1: Hyperparameters for Baselines and InferenceGuard.
These hyperparameter settings ensure a fair and controlled comparison of the baseline methods and the proposed InferenceGuard model. We compare the performance of baseline methods, with reward model, cost model and lagriangian multiper setting, under these hyperparameter settings, and summarised the result in Table 2.
When sampling from the base model, we used a temperature of 1.0 without top_k nor top_p.
Table 2: Performance Comparison on Dataset PKU-SafeRLHF
Method Average Reward Average Cost Safety Rate Inference Time (s)
Alpaca-7B Base 6.15 ($\pm$ 1.51) 1.33 29.47% 1.1
RECONTROL 6.2 ($\pm$ 1.56) 1.33 29.5% 1.7
RECONTROL + Lagrangian multiplier 6.19 ($\pm$ 1.50) 1.33 29.7% 2
Best-of-N + Lagrangian multiplier $N=100$, $\lambda=5$ 5.35 ($\pm$ 1.62) -0.46 48.22% 29
Best-of-N + Lagrangian multiplier $N=200$, $\lambda=5$ 5.25 ($\pm$ 1.64) -0.72 54.2% 58
Best-of-N + Lagrangian multiplier $N=500$, $\lambda=5$ 6.04 ($\pm$ 1.85) -1.27 52.17% 145
Best-of-N + Lagrangian multiplier $N=500$, $\lambda=10$ 5.51 ($\pm$ 1.66) -1.44 54.01% 145
Best-of-N + Augmented safety $N=200$ 7.51 ($\pm$ 1.89) 0.67 60.07% 58
Best-of-N + Augmented safety $N=500$ 7.78 ($\pm$ 2.09) 0.42 65.74 % 145
Beam search + Lagrangian multiplier $N=128$, $\lambda=5$ 6.58 ($\pm$ 1.95) -1.02 50.19% 32
Beam search + Lagrangian multiplier $N=256$, $\lambda=5$ 6.69 ($\pm$ 2.08) -1.28 52.43% 60
Beam search + Augmented safety $N=128$ 8.29 ($\pm$ 2.02) 0.64 58.89% 39
Beam search + Augmented safety $N=256$ 8.69 ($\pm$ 2.15) 0.55 61.79 % 82
ARGS $\omega=2.5$ 6.74 ($\pm$ 1.70) 1.47 28.19% 82
ARGS + Lagrangian multiplier $\omega=2.5$ 3.21 ($\pm$ 1.59) -0.85 75.8% 111
ARGS + Cost Model $\omega=2.5$ 0.19 ($\pm$ 1.65) -2.21 81.6% 78
InferenceGuard (N=128) 7.08 ($\pm$ 2.49) -0.63 88.14% 65
[…]
Beaver-7B-v3 Base 5.83 ($\pm$ 1.62) -2.89 75.89% 1.2
RECONTROL 5.9 ($\pm$ 1.56) -2.90 75.9% 2
RECONTROL + Lagrangian multiplier 5.91 ($\pm$ 1.50) -2.91 75.9% 2.6
Best-of-N + Lagrangian multiplier $N=100$ 6.52 ($\pm$ 1.88) -3.63 85.7% 40
Best-of-N + Lagrangian multiplier $N=200$ 6.61 ($\pm$ 1.89) -3.62 85.8% 58
Best-of-N + Augmented safety $N=100$ 8.55 ($\pm$ 1.58) -2.96 97.23% 40
Best-of-N + Augmented safety $N=200$ 9.01 ($\pm$ 1.63) -2.98 97.76% 58
Beam search + Lagrangian multiplier $N=64$, $\lambda=5$ 8.33 ($\pm$ 1.79) -4.09 87.08% 36
Beam search + Lagrangian multiplier $N=128$, $\lambda=5$ 8.63 ($\pm$ 1.80) -4.21 87.35 % 64
Beam search + Augmented safety $N=64$ 9.84 ($\pm$ 1.4) -2.93 95.38 % 22
Beam search + Augmented safety $N=128$ 10.31 ($\pm$ 1.37) -2.94 97.36% 39
ARGS $\omega=2.5$ 6.72 ($\pm$ 1.83) -2.59 78.5% 94
ARGS $\omega=2.5$ + Lagrangian multiplier 2.26 ($\pm$ 1.56) -1.64 81% 127
ARGS $\omega=2.5$ + Cost Model 0.01 ($\pm$ 1.37) -3.27 98.4% 90
InferenceGuard (N=128) 10.26 ($\pm$ 1.42) -2.96 99.7% 39
InferenceGuard with Critic (N=128) 10.27 ($\pm$ 1.50) -2.94 100% 39
D.3 Critic Network and Training Process
This section outlines the critic network architecture used for InferenceGuard. The critic network is designed to estimate the cost of partial responses and guide optimization during inference. We assume that trajectories terminate at a maximum time $T$, and the critic aims to predict the sign of the safety compliance metric $\text{z}_{T}$, and the discounted cumulative task cost $\gamma^{T}\bar{c}_{\text{task}}$.
Critic Network Architecture
The critic network takes two types of input: the hidden states ($\bm{h}_{t}$) and the key-value pairs ($\mathbf{o}_{t}$), representing contextual and state information, respectively. These are passed through a series of layers to estimate the required outputs. The network utilizes downscaling and attention layers to reduce the dimensionality of the input data, ensuring efficient processing of large-scale representations.
In terms of model size, the total parameter count of the critic network is approximately 0.7 billion parameters, providing a balance between model capacity and computational efficiency.
Training Process
The critic network is trained using a combination of optimization techniques aimed at predicting the safety compliance and task cost accurately. The network is optimized with the hyperparameters in Table 3.
Hyperparameter Value
Hidden Dimension 4096
Learning Rate $1\times 10^{-5}$
Number of Epochs 50
Discount Factor ($\gamma$) 0.999
Batch Size 8
Safety Budget $d$ 10
Table 3: Hyperparameters for Critic Network Training.
During training, the network is fed batches of hidden states and key-value pairs, and the weights are updated to minimize the loss between predicted and true values. The critic network’s ability to predict both the safety compliance and task cost ensures it can guide the optimization process at inference time, while adhering to safety constraints.
The model uses a penalty term to enforce the safety budget constraint. This penalty discourages the network from violating the safety threshold, steering for safer responses during intervention.
D.4 Qualitative Comparisons
We present several examples from the PKU-SafeRLHF test dataset to demonstrate how InferenceGuard steers the base model outputs towards safer responses. These examples highlight differences in response safety between InferenceGuard and baseline methods, including BoN, Beam Search, RECONTROL, and ARGS, evaluated on the Alpaca and Beaver-v3 models, as shown in Figure 4 and 5. In each case, InferenceGuard successfully adheres to safety constraints while maintaining task performance.
(a) Alpaca Example 1
(b) Alpaca Example 2
Figure 4: Generated response by different methods on the Alpaca-7B
(a) Beaver Example 1
(b) Beaver Example 2
Figure 5: Generated response by different methods on the Beaver-v3-7B