### v3

# On Almost Surely Safe Alignment of Large Language Models at Inference-Time

Abstract
We introduce a novel inference-time alignment approach for LLMs that aims to generate safe responses almost surely, i.e., with probability approaching one. Our approach models the generation of safe responses as a constrained Markov Decision Process (MDP) within the LLM’s latent space. We augment a safety state that tracks the evolution of safety constraints and dynamically penalize unsafe generations to ensure the generation of safe responses. Consequently, we demonstrate formal safety guarantees w.r.t. the given cost model upon solving the MDP in the latent space with sufficiently large penalties. Building on this foundation, we propose InferenceGuard, a practical implementation that safely aligns LLMs without modifying the model weights. Empirically, we demonstrate that InferenceGuard effectively balances safety and task performance, outperforming existing inference-time alignment methods in generating safe and aligned responses. Our findings contribute to the advancement of safer LLM deployment through alignment at inference-time, thus presenting a promising alternative to resource-intensive, overfitting-prone alignment techniques like RLHF.
Contains potentially harmful examples.
1 Introduction
LLMs have demonstrated impressive capabilities across a diverse set of tasks, such as summarization [1, 2], code generation [3, 4], and embodied robotics [5, 6]. However, since those models are primarily trained on vast, unsupervised datasets, their responses can often be biased, inaccurate, or harmful [7, 8, 9, 10]. To prevent such controversial content, LLMs require alignment with better human values.
The predominant approach for LLM alignment is Reinforcement Learning from Human Feedback (RLHF) [11, 12], which fine-tunes the model using human preference data. A drawback of RLHF, however, is its potential training cost and risk of overfitting, partly because this method modifies the model’s weights. In contrast, inference-time alignment adjusts the model’s outputs directly during inference to align with a reward model, while leaving the model weights fixed ([13, 2, 14]). Despite the successes of inference-time alignment, their safety aspects have received limited attention so far.
In this work, we aim to develop a principled inference-time alignment technique that guarantees the safety of LLM responses almost surely, i.e., with a probability approaching one. To do so, we reformulate the safe generation of inference-time responses as an instance of constrained Markov decision processes (cMDP) with the objective of maximizing task performance while satisfying safety cost constraints. We map the cMDP to an unconstrained one through safety state augmentation, bypassing Lagrangian approaches’ limitations that struggle to balance reward maximization and safety feasibility. Specifically, the tracking of the safety state enables one to dynamically penalize the task reward when constraint violations occur. Focusing on practical efficiency, we adopt a critic-based approach to solve the augmented MDP, eliminating the need for gradients in the LLM. To ensure efficiency, we train our critic in the LLM’s latent space, keeping it small in size and fast during inference. This shift to the latent space complicates the theoretical framework, requiring extensions from previous works [15, 16]. By doing so, we establish, for the first time, that for sufficiently large penalties one can guarantee almost sure safety in the original token space w.r.t. given cost model.
Figure 1: Overview of the InferenceGuard framework. Given a prompt $\mathbf{x}$, InferenceGuard sequentially generates beams of tokens from the base LLM, augments them with the safety state to track the evolution of safety constraints, evaluates each beam using our learned value models for both safety and task alignment, and filters the top $K$ beams (see Section 5.2). If all beams are unsafe, it penalizes the logits of unsafe tokens and resamples. Consequently, InferenceGuard efficiently generates responses with superior safety rates and strong task performance.
To leverage this theoretical guarantee in practice, we build upon the augmented MDP framework and introduce two novel implementations for safe inference-time alignment: one that learns a compact critic in the latent space for cases where safety costs can only be queried after the generation of complete responses and another that leverages direct cost queries for efficient inference-time optimization. Finally, we integrate these components into a lookahead algorithm (e.g., Beam Search or Blockwise Decoding [14]) proposing InferenceGuard. While test-time alignment often introduces additional latency – commonly referred to as the alignment tax – InferenceGuard matches the decoding cost of standard beam search yet achieves markedly high safety rates – 94.46% on Alpaca-7B, 98.45% on Llama3.1-8B-Instruct, 98.97% on Vicuna-7B and 100% on Beaver-7B-v3. Notably, this was accomplished while maintaining a strong balance with rewards, setting new state-of-the-art.
2 Related Work
LLM alignment and safety: Pre-trained LLMs are often aligned to specific tasks using RLHF, where LLMs are fine-tuned with a learned reward model [12, 2, 17, 11] constructed from human feedback using standard reinforcement learning algorithms like PPO [18]. More recent approaches, such as [19, 20, 21, 22, 23, 24, 25, 26], bypass reward learning and instead align pre-trained models directly with human preferences. [27, 8] first applied fine-tuning in the context of safety, and [28] proposed safe fine-tuning via Lagrangian optimization. Other safety-focused methods such as [29, 30, 31, 32, 33, 34, 35, 36, 37] are either orthogonal, handle a different problem to ours, or can not ensure almost sure safety w.r.t. the given cost model during inference.
Inference time alignment: In order to reduce reliance on resource-intensive and often hard-to-stabilize RL processes inherent in the RLHF paradigm, inference-time alignment techniques such as Best-of-N [13, 2, 38, 39], guided decoding (steering token generation based on reward or a trained value function) [40, 41, 14, 42, 43, 44, 45, 46], among others, have been proposed. Importantly, these techniques are designed modularly, allowing the alignment module to integrate seamlessly with the pre-trained model by adjusting the model’s responses directly at inference time. This modularity enables flexible inference-time reconfigurability and quick adaptation to new reward models and datasets.
While a few works have attempted to tackle the safety issue in inference-time-aligned responses, they mainly focus on prompt-based alignment [32, 33, 47, 48], trainable safety classifiers [49, 50] or protections against adversarial attacks and jailbreaks [51, 34, 52, 53]. That said, prompt-based methods cannot be guaranteed to consistently produce safe responses, as ensuring safety is heavily reliant on user intervention, requiring extensive engineering and expertise to manage the model’s output effectively. Trainable classifiers focus only on the safety of decoded responses using hidden states or virtual tokens, ignoring task alignment and lacking theoretical guarantees. Moreover, while adversarial robustness is crucial, our work focuses on the key challenge of generating inherently safe responses from the LLM. Compared to those methods, we are the first to theoretically guarantee almost surely safe alignment w.r.t. given cost model with strong empirical results. Operating in the latent space enables us to train smaller, inference-efficient critics while optimally balancing rewards and safety constraints without introducing extra parameters, e.g., Lagrangian multipliers.
We detail other related works extensively in Appendix A.
3 Background
LLMs can be viewed as stochastic dynamical systems, where the model’s behavior evolves probabilistically over time, governed by its internal parameters and the inputs it receives. In this perspective, each new token is generated based on the model’s evolving hidden state [42, 54]. Formally, an LLM transitions as follows: $\left[\mathbf{h}_{t+1},\mathbf{o}_{t+1}\right]^{\mathsf{T}}=f_{\text{LLM}}(\mathbf{h}_{t},\text{y}_{t})$, with $\text{y}_{t}\sim\text{SoftMax}(\mathbf{W}\mathbf{o}_{t})$. Here, $f_{\text{LLM}}(\cdot)$, denotes the aggregation of all decoding layers, $\text{y}_{t}$ a generated token at each time step $t$, and $\mathbf{o}_{t}$ the logits which are linearly mapped by $\mathbf{W}$ to produce a probability distribution over the vocabulary space. Moreover, $\mathbf{h}_{t}$ comprises all key-value pairs accumulated from previous time steps  [^1] Note, $\mathbf{h}_{t}=\left\{\mathbf{K}^{(l)}_{j},\mathbf{V}_{j}^{(l)}\right\}_{l=1}^{L}$ for $j\in[1,t]$, i.e., keys and values from all layers till time step $t$.. The system evolves until the end-of-sequence (EOS) token is reached.
Test-Time Alignment of LLMs. A pre-trained LLM can be ensured to generate outputs consistent with desired behaviors by solving an MDP, whose initial state is determined by the test prompt. Rewards/costs for alignment can come from various sources, such as human feedback [19, 55], environmental feedback from the task or verifiers [56, 57, 58, 59, 60, 5], or pre-trained reward models [61, 62, 63].
Several approaches address the challenge of test-time alignment. For instance, beam search with rewards [64] extends traditional beam search by integrating a reward signal to guide LLM decoding at test time. Monte Carlo tree search, on the other hand, takes a more exploratory approach by simulating potential future token sequences to find the path that maximizes the reward function [62]. Best-of-N (BoN) generates multiple candidate sequences and selects the one with the highest reward  [2, 13, 38]. We focus on beam search methods for more scalability when developing InferenceGuard.
4 Safe Test-Time Alignment of LLMs
We frame the problem of safely aligning LLMs at test time as a constrained Markov decision process (cMDP). As noted in [65, 16] a cMDP is defined as the following tuple: $\mathcal{M}=\left\langle\mathcal{S},\mathcal{A},\mathcal{C}_{\text{task}},\mathcal{C}_{\text{safety}},\mathcal{P},\gamma\right\rangle$, with $\mathcal{S}$ and $\mathcal{A}$ denoting the state and action spaces, respectively. The cost function $\mathcal{C}_{\text{task}}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ dictates the task’s cost [^2] We define costs as negative rewards, which transforms the problem into an equivalent cost-based MDP., while $\mathcal{C}_{\text{safety}}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ represents a safety cost, which encodes the constraints that the actor must satisfy during inference. The transition model $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ captures the probability of transitioning to a new state given the current state and action. Meanwhile, the discount factor $\gamma\in[0,1)$ trades off immediate versus long-term rewards. The goal of constrained MDPs is to find a policy $\pi:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ that minimizes the task’s cost while simultaneously satisfying the safety constraints. Given a safety budget d, we write:
$\min_{\pi}$ $\mathbb{E}_{\mathcal{P},\pi}\Big{[}\sum_{t}\gamma^{t}\mathcal{C}_{\text{task}}(\mathbf{s}_{t},\mathbf{a}_{t})\Big{]}\quad\text{s.t.}\quad\mathbb{E}_{\mathcal{P},\pi}\Big{[}\sum_{t}\gamma^{t}\mathcal{C}_{\text{safety}}(\mathbf{s}_{t},\mathbf{a}_{t})\Big{]}\leq d,$ (1)
4.1 Safe Test-Time Alignment as cMDPs
We treat the generation process of safe test-time alignment of LLMs as the solution to a specific cMDP. We introduce our state variable $\mathbf{s}_{t}=\{\mathbf{x},\mathbf{y}_{<t}\}$, which combines the input prompt $\mathbf{x}$ with the tokens (or partial responses) decoded until step $t$. Our policy generates a new token $\text{y}_{t}$ that we treat as an action in the model’s decision-making process. The transition function $\mathcal{P}$ of our MDP is deterministic, where the state $\mathbf{s}_{t}$ is updated by incorporating the generated action $\text{y}_{t}$, i.e., $\mathbf{s}_{t+1}=\mathbf{s}_{t}\oplus\text{y}_{t}=\{\mathbf{x},\mathbf{y}_{\leq t}\}$. We also assume the existence of two cost functions $\mathcal{C}_{\text{task}}$ and $\mathcal{C}_{\text{safety}}$ to assess the correctness and safety of the LLM’s responses. As described in Section 3, those functions can originate from various sources, such as human feedback, environmental verifiers, or pre-trained models. While we conduct experiments with these functions being LLMs (see Section 6), our method can be equally applied across various types of task and safety signals.
We assume the availability of a function $\mathcal{C}_{\text{task}}$ that evaluates the alignment of the LLM with the given task. This function assigns costs to the partial response based on the input prompt $\mathbf{x}$ such that:
$\mathcal{C}_{\text{task}}([\mathbf{x},\mathbf{y}_{\leq t}]):=\begin{cases}0&\text{if }\text{y}_{t}\neq\texttt{EOS}\\ c_{\text{task}}([\mathbf{x},\mathbf{y}_{\leq t}])&\text{if }\text{y}_{t}=\texttt{EOS}\end{cases}$ (2)
For the safety cost $\mathcal{C}_{\text{safety}}$, we assume the function assigns non-zero costs to any partial answer without waiting for the final token. This is crucial because we want to flag unsafe responses early rather than waiting until the end of the generation process. Many pre-trained models are available for this purpose on Hugging Face, which we can leverage—more details can be found in Section 6. With this, we write safe test-time alignment as an instance of Equation 1:
$\min_{\pi}$ $\mathbb{E}_{\pi}\Big{[}\sum_{t}\gamma^{t}\mathcal{C}_{\text{task}}(\overbrace{\{\mathbf{x},\mathbf{y}_{<t}\}}^{\mathbf{s}_{t}},\text{y}_{t})\Big{]}\quad\text{s.t.}\quad\mathbb{E}_{\pi}\Big{[}\sum_{t}\gamma^{t}\mathcal{C}_{\text{safe}}(\{\mathbf{x},\mathbf{y}_{<t}\},\text{y}_{t})\Big{]}\leq d.$ (3)
The above objective aims to minimize the task cost $\mathcal{C}_{\text{task}}$ while ensuring that the safety cost does not exceed a predefined budget $d$. The expectation is taken over the actions (tokens) generated at each step.
4.2 State Augmented Safe Inference-Time Alignment
We could technically use off-the-shelf algorithms to solve Equation 3, such as applying a Lagrangian approach as proposed in [28]. However, there are two main issues with using these standard algorithms. First, they generally require gradients in the model itself—specifically, the LLM—which we want to avoid since our goal is to perform inference-time alignment without retraining the model. Second, these methods rely on a tunable Lagrangian multiplier, making it challenging to maximize rewards while satisfying almost sure constraints optimally.
Instead of a Lagrangian approach, we take a different direction by augmenting the state space and extending the method proposed by [16] to large language models. In our approach, we augment the state space of the constrained MDP with a “constraint tracker”, effectively transforming the problem into an unconstrained one. This allows us to apply Bellman equations and conduct rigorous proofs with almost sure constraint satisfaction results. However, applying the techniques and proofs from [16] to our test-time setting is not entirely straightforward due to two main challenges: first, the differences in the constrained MDP setting, and second, the process by which we train critics, as we will demonstrate next.
Augmenting the State Space. The following exposition builds on [16], extending their method to address LLM-specific challengers, an area they did not cover. The core idea is to transform the constrained MDP into an unconstrained one by augmenting the state with an additional variable that tracks the remaining budget of the constraint. While doing so, we must ensure that: PI) our augmented state variable tracks the constraints and maintains the Markovian nature of transition dynamics; and PII) our task cost $\mathcal{C}_{\text{task}}$ accounts for this new state representation.
We solve PI by tracking a scaled-version of the remaining safety budget (see Equation 3 and) $\mathbf{\omega}_{t}=d-\sum_{k=1}^{t}\gamma^{k}\mathcal{C}_{\text{safety}}(\{\mathbf{x},\mathbf{y}_{<k}\},\text{y}_{k})$, defined as $\text{z}_{t}=\mathbf{\omega}_{t-1}/\gamma^{t}$. The update of $\text{z}_{t}$ satisfies:
$\text{z}_{t+1}=(\mathbf{\omega}_{t-1}-\gamma^{t}\mathcal{C}_{\text{safety}}(\{\mathbf{x},\mathbf{y}_{<t}\},\text{y}_{t}))/\gamma^{t+1}=(\text{z}_{t}-\mathcal{C}_{\text{safety}}(\{\mathbf{x},\mathbf{y}_{<t}\},\text{y}_{t}))/\gamma,\ \ \text{with $\text{z}_{0}=d$.}$ (4)
The dynamics of $\text{z}_{t}$ are Markovian and dependent only on $\text{z}_{t-1}$, $\text{y}_{t-1}$ and current state $\{\mathbf{x},\mathbf{y}_{<t-1}\}$. Hence, we can easily augment our original state space with $\text{z}_{t}$, such that $\tilde{\mathbf{s}}_{t}=[\mathbf{s}_{t},\text{z}_{t}]=[\{\mathbf{x},\mathbf{y}_{<t}\},\text{z}_{t}]$. The original dynamics can also be redefined to accommodate for $\tilde{\mathbf{s}}_{t}$:
[…]
Concerning PII, we note that enforcing the original constraint in Equation 3 is equivalent to enforcing an infinite number of the following constraints:
$\sum_{k=0}^{t}\gamma^{k}\mathcal{C}_{\text{safety}}(\{\mathbf{x},\mathbf{y}_{<k}\},\text{y}_{k})\leq d\ \ \forall t\geq 1.$ (5)
As noted in [16], this observation holds when the instantaneous costs are nonnegative, ensuring that the accumulated safety cost cannot decrease. In our case, it is natural to assume that the costs are nonnegative for LLMs, as safety violations or misalignments in the output typically incur a penalty, reflecting the negative impact on the model’s performance or ethical standards.
Clearly, if we enforce $\text{z}_{t}\geq 0$ for all $t\geq 0$, we automatically get that $\mathbf{\omega}_{t}=d-\sum_{k=0}^{t}\gamma^{k}\mathcal{C}_{\text{safety}}(\{\mathbf{x},\mathbf{y}_{<k}\},\text{y}_{k})\geq 0$ for all $t\geq 0$, thus satisfying the infinite constraints in Equation 5. We can do so by reshaping the tasks’s instantaneous cost to account for the safety constraints:
$\tilde{\mathcal{C}}_{\text{task}}^{\infty}(\tilde{\mathbf{s}}_{t},\text{y}_{t}):=\begin{cases}\mathcal{C}_{\text{task}}([\mathbf{x},\mathbf{y_{\leq t}}])&\text{z}_{t}>0\\ +\infty&\ \text{z}_{t}\leq 0,\end{cases}$ (6)
[…]
$\min_{\pi}\mathbb{E}_{\pi}\Big{[}\sum_{t}\gamma^{t}\tilde{\mathcal{C}}_{\text{task}}^{\infty}(\tilde{\mathbf{s}}_{t},\text{y}_{t})\Big{]}.$ (7)
Using gradient-based techniques, one could optimize the augmented MDP in Equation 7. However, since our goal is to enable safety at test time without retraining, we adopt a critic-based approach that does not require gradients during inference, as we show next.
5 InferenceGuard: Safety at Test-Time
When designing our critic, we considered several crucial factors for test-time inference. These included its size, ease of training for quick adaptation, and flexibility to operate in real-time without significant latency. As such, we chose to train the critic in the latent space of the LLM rather than directly in the textual space, enabling a more efficient solution that meets the constraints of test-time alignment.
[…]
To define our critic, we require a representation of our augmented state $\tilde{\mathbf{s}}_{t}=[\mathbf{s}_{t},\text{z}_{t}]$ within the latent space. As noted above, we can acquire $(\mathbf{h}_{t},\mathbf{o}_{t})$ from the transformer architecture. We call this mapping $\phi$, whereby $(\mathbf{h}_{t},\mathbf{o}_{t})=\phi(\{\mathbf{x},\mathbf{y}_{<t}\})$. To embed $\text{z}_{t}$, we use an identity mapping which enables us to input the actual tracking of the constraints directly to the critic without any loss of information.
5.1 Theoretical Insights
In this section, we show that optimizing in the latent space preserves safety constraints in the original token space and prove that our approach guarantees almost sure safety w.r.t. given safety cost model.
We consider two essential questions: i) Can we compute an optimal policy in the latent space?, ii) If we enforce safety constraints in the latent space, do they still hold in the original token space? While [16] established theoretical results for safety-augmented MDPs in standard (non-LLM) RL settings, their work does not address how guarantees in the latent space translate to the original token space. To handle those problems, we extend the theorems from [15, 16] to ensure the following properties:
  • Prop I) The latent MDP indeed satisfies the Bellman equations (Theorem 5.1 (a)) and, hence, allows us to compute an optimal policy in this space,
  • Prop II) The latent space policies and value functions are valid in the original token space. Hence, optimizing in the latent space preserves the constraints in original token space (Theorem 5.1 (b,c))
  • Prop III) The resulting policy satisfies safety constraints almost surely (Theorem 5.1), meaning if a policy is safe in the latent and original token space with finite expected cost w.r.t. Equation 7, it is also almost surely safe in the actual LLM token space.
We begin by defining the latent space MDP’s cost and transition function:
Definition 5.1.
$\exists\phi(\cdot)$ and functions $\bar{\mathcal{C}}^{n}_{\text{task}}$ and $\bar{\mathcal{P}}$ such that:
[…]
$\bar{\mathcal{P}}(\phi(\{\mathbf{x},\mathbf{y}_{\leq t}\}),\text{z}_{t+1}|\phi(\{\mathbf{x},\mathbf{y}_{<t}\}),\text{z}_{t},\text{y}_{t})={\mathcal{P}}(\tilde{\mathbf{s}}_{t+1}|\tilde{\mathbf{s}}_{t},\text{y}_{t}),$
where $\tilde{\mathbf{s}}_{t}$ is the augmented state in the original token space. Definition 5.1 ensures that the cost incurred by the augmented state $\tilde{\mathbf{s}}_{t}=[\{\mathbf{x},\mathbf{y}_{<t}\},\text{z}_{t}]$ w.r.t. $\tilde{\mathcal{C}}^{n}_{\text{task}}$ is equal to the latent cost incurred by the latent state $[\phi(\{\mathbf{x},\mathbf{y}_{<t}\}),\text{z}_{t}]$ w.r.t $\bar{\mathcal{C}}_{\text{task}}^{n}$. Moreover, it ensures that the transition dynamics of the augmented state in the original token space and the corresponding latent state in the latent space are equivalent.
This equivalence enables us to derive an optimal policy for the latent MDP and apply it to minimize the cost objective in the original augmented MDP (see Equation 7). We proceed to analyze the existence of such an optimal policy in the latent space through the following standard assumptions [16] on $\bar{\mathcal{C}}^{n}_{\text{task}}$, and $\bar{\mathcal{P}}$: A1. The function $\bar{\mathcal{C}}_{\text{task}}^{n}(\mathbf{h},\mathbf{o},\text{z},\text{y})$ is bounded, measurable, nonnegative, lower semi-continuous w.r.t. $(\mathbf{h},\mathbf{o},\text{z})$ for a given y, and A2. The transition law $\bar{{\mathcal{P}}}$ is weakly continuous for any y.
Next, we define $\bar{\pi}$ as a policy in the latent space that maps $(\mathbf{h},\mathbf{o},\text{z})\rightarrow\text{y}$, and its value function for an initial state $(\mathbf{h}_{0},\mathbf{o}_{0},\text{z}_{0})$ as follows: $\bar{V}^{n}(\bar{\pi},\mathbf{h}_{0},\mathbf{o}_{0},\text{z}_{0})=\mathbb{E}_{\bar{\pi}}\Big{[}\sum_{t=0}^{\infty}\gamma^{t}\bar{\mathcal{C}}_{\text{task}}^{n}(\mathbf{h}_{t},\mathbf{o}_{t},\text{z}_{t},\text{y})\Big{]}$. Then, one can define the optimal value function:
$\bar{V}^{\star,n}(\mathbf{h},\mathbf{o},\text{z})=\min_{\bar{\pi}}\bar{V}^{n}(\bar{\pi},\mathbf{h},\mathbf{o},\text{z}).$ (8)
Since we cannot optimize directly in the original constrained MDP, we first show that solving for an optimal policy in the latent MDP preserves key properties of the original problem. The following theorem formalizes this by proving the existence of the optimal policy and its mapping to the original MDP.
{restatable}
thmthmsauteequivalence(Optimality in the Latent Space) Given A1-A2, the latent MDP in Definition 5.1 satisfies:
  1. a)
    (Prop I) For any finite $n$, the Bellman equation holds, i.e., there exists $\bar{V}^{\star,n}(\mathbf{h},\mathbf{o},\text{z})$ such that:
    $\bar{V}^{\star,n}(\mathbf{h},\mathbf{o},\text{z})=\min_{\text{y}\in{\mathcal{V}}}\Big{(}\bar{\mathcal{C}}^{n}_{\text{task}}(\mathbf{h},\mathbf{o},\text{z},\text{y})+\gamma\bar{V}^{\star,n}(\mathbf{h}^{\prime},\mathbf{o}^{\prime},\text{z}^{\prime})\Big{)},(\mathbf{h}^{\prime},\mathbf{o}^{\prime},\text{z}^{\prime})\sim\bar{{\mathcal{P}}}(\cdot|\mathbf{h},\mathbf{o},\text{z},\text{y})$
    Furthermore, the optimal policy solving Equation 8 has the representation $y\sim\bar{\pi}^{\star,n}(\cdot\mid\mathbf{h},\mathbf{o},\text{z})$;
  2. b)
    (Prop II) The optimal value functions $\bar{V}^{\star,n}$ converge monotonically to $\bar{V}^{\star,\infty}$.
  3. c)
    (Prop II) The optimal policy in the latent space $\bar{\pi}^{\star,n}$ is also optimal in the original token space if used as $\bar{\pi}^{\star,n}(\phi(\cdot))$, minimizing Equation 7, even as $n\rightarrow\infty$.
The above theorem ensures that finding and securing the existence of the optimal policy in the latent space is sufficient to solve Equation 7 optimally. Informally, the latent space acts as a faithful representation, preserving constraints and making optimization computationally efficient. This implies that the optimal policies and value functions in the latent space remain valid in the original space. We relegate the proof to Appendix B.3 but outline the main steps for Prop II below:
We first show the infimum operator recovers lower semi continuity, which implies for finite $n$ ,$\bar{V}^{\star,n}$ is lower semicontinous due to assumptions A1, A2. Next, we formulate $\bar{V}^{\star,n}\to\bar{V}^{\star,\infty}$, as the convergence of limit infimum of a sequence of increasing lower semi continuous functions, $v_{n}\to v_{0}$, to infimum of $v_{0}$. We construct a decreasing sequence of compact action sets consisting of the infimum actions $a_{n}$ of $v_{n}$ and converging to the infimum action set of $v_{0}$ consisting of $a_{0}$. We use properties of the discrete action space to show that this sequence of actions has a convergent subsequence $a_{n_{i}}\to a_{0}$ and use that to show that limit infimum of $v_{n}$ converges to infimum of $v_{0}$ (see part-c of Lemma 1 ).
Now, we derive Prop III that ensures the safety cost constraints are almost surely satisfied. This is more challenging than Equation 3, where only the expected safety cost is constrained:
[…]
While the formulation in Equation 9 is “stronger” than Equation 3, solving for the augmented MDP formulation with objective as Equation 7 can yield a policy satisfying the above almost sure constraints. We formally state this result in Theorem 5.1 and relegate the proof to Appendix B.3. {restatable}thmthmalmostsure (Almost Sure Safety) Consider an augmented MDP with cost function $\tilde{\mathcal{C}}_{\text{task}}^{\infty}$. Suppose an optimal policy exists $\pi^{\star}$ solving Equation 7 (see Theorem 5.1) with a finite cost, then $\pi^{\star}$ is an optimal policy for Equation 9, i.e., $\pi^{\star}$ is safe with probability approaching one or almost surely.
Theorem 5.1 justifies our state-augmented MDP reformulation, as it guarantees that a solution to the augmented MDP (Equation 7) almost surely satisfies the constraints of Equation 3 distinguishing its effectiveness from alternative reformulations that may not offer such a guarantee.
5.2 Algorithm and Practical Implementation
Building on our theoretical framework, we propose a search algorithm with two approaches: i) pre-training a small latent-space critic for cases where costs are available only for complete trajectories and ii) directly leveraging intermediate costs for search optimization.
Training a Latent-Space Critic. We make the usual assumption that trajectories terminate at a maximum length $T$. In this case, the value function simplifies to become: $\bar{V}^{n}(\mathbf{h}_{t},\mathbf{o}_{t},\text{z}_{t})=\mathbb{E}_{\bar{\pi}}[\gamma^{T}\bar{c}_{\text{task}}({\mathbf{h}}_{T},{\mathbf{o}}_{T})]$ if there is safety budget left, i.e., if $\text{z}_{T}>0$, or $n$ if $\text{z}_{T}\leq 0$, where $\bar{c}_{\text{task}}({\mathbf{h}}_{t},{\mathbf{o}}_{t})=c_{\text{task}}([\mathbf{x},\mathbf{y}_{\leq t}])$ in the latent MDP. Hence, it is sufficient to predict: the sign of $\text{z}_{T}$ and the value of $\gamma^{T}\bar{c}_{\text{task}}({\mathbf{h}}_{T},{\mathbf{o}}_{T})$ to assess the quality of a state. We estimate those through Monte Carlo (MC) sampling. Specifically, we generate multiple trajectories from the initial state ($\mathbf{h}_{0},\mathbf{o}_{0},\text{z}_{0})$ using the reference policy, and compute the mean terminal cost, the sign of $z_{T}$ to serve as targets for the critic training. The usual alternative to MC sampling is Temporal Difference (TD) learning, where the critic is updated based on the difference between the current estimate and a bootstrapped estimate from the next state. However, MC sampling offers two advantages: i) it simplifies training by using separate supervised signals for quality and safety, unlike TD, which combines both, and ii) it allows dynamic adjustment of $n$ without retraining.
We train a critic network with two heads ( $f^{1}_{\mathbf{\theta}}$ and $f^{2}_{\mathbf{\theta}}$) by sampling responses from the base model and scoring them using the cost function. We define $\mathcal{J}_{1}$ as the binary cross-entropy for predicting the sign of $\text{z}_{T}$ and $\mathcal{J}_{2}$ as the mean squared error for predicting $\gamma^{T}\bar{c}_{\text{task}}({\mathbf{h}}_{T},{\mathbf{o}}_{T})$. Our critic training minimizes: $\mathcal{J}(\mathbf{\theta})=\mathbb{E}_{\bar{\pi}}\Big{[}\sum_{t=1}^{T}\mathcal{J}_{1}\left(f^{1}_{\mathbf{\theta}}({\mathbf{h}}_{t},{\mathbf{o}}_{t},\text{z}_{t}),\text{z}_{T}>0\right)+\mathcal{J}_{2}\left(f^{2}_{\mathbf{\theta}}({\mathbf{h}}_{t},{\mathbf{o}}_{t},\text{z}_{t}),\gamma^{T}\bar{c}_{\text{task}}({\mathbf{h}}_{T},{\mathbf{o}}_{T})\right)\Big{]}$.
Search method. We build on the beam search strategy from [14, 66] wherein we sequentially sample $N$ beams of $d$ tokens from the pre-trained model and choose $K$ beams with the highest scores as possible continuations of the prompt (see Algorithm 1 in Appendix C). This ensures that we focus on the most promising continuations. The goal of the scoring function is to balance the immediate task cost and the predicted future task cost while ensuring safety. This is repeated until we complete trajectories. Given a token trajectory $\text{y}_{t:t+d}$, we present a scoring function $\text{E}_{\text{critic}}$ that assumes we cannot evaluate intermediate answers with the cost functions. However, when immediate safety costs are available, a simpler scoring function can be used, see Appendix C. We define $\text{E}_{\text{critic}}$ as:
$\text{E}_{\text{critic}}(\text{y}_{t:t+d})=$$\begin{cases}\gamma^{T}\bar{c}_{\text{task}}(\cdot)&t+d=T\text{ and }\text{z}_{t+d}>0\\ n&t+d=T\text{ and }\text{z}_{t+d}\leq 0\\ f^{2}_{\mathbf{\theta}}(\cdot)&f^{1}_{\mathbf{\theta}}(\cdot)>0.5\\ n&\text{otherwise}.\end{cases}$
This $\text{E}_{\text{critic}}$ scoring function evaluates token sequences by balancing safety and task performance. At the final step ($t+d=T$), it assigns a score based on the task cost $\mathcal{C}_{\text{task}}$ if safety constraints are met ($\text{z}_{t+d}>0$); otherwise, it applies a high penalty $n$. For intermediate steps, it relies on a trained critic. If the critic confidently predicts safety ($f_{\mathbf{\theta}}^{1}(\mathbf{h}_{t+d},\mathbf{o}_{t+d},\text{z}_{t+d})>0.5$), it uses the estimated future cost ($f_{\mathbf{\theta}}^{2}(\mathbf{h}_{t+d},\mathbf{o}_{t+d},\text{z}_{t+d})$); otherwise, it assigns the penalty $n$ as a conservative safeguard.
Figure 2: Trade-offs between safety, reward, and inference time evaluated on Alpaca-7B and Beaver-v3-7B using the PKU-SafeRLHF dataset (top), and Vicuna-7B and LLaMA3-8B using the HEx-PHI and HH-RLHF datasets respectively (bottom). Each subplot visualizes the average reward score versus safety rate, and the inference time denoted by the color of each symbol. InferenceGuard achieves a well-balanced trade-off across all three objectives by positioning in the Optimal Region.
Sampling Diversity. Finally, if the right selection strategy can guarantee that we will converge on a safe solution, it does not consider how many samples would be necessary. To increase the search speed, we introduce a diversity term in the sampling distribution when no safe samples were found based on the token frequency of failed beams. We denote $F$ as the frequency matrix counting the tokens we previously sampled from $t$ to $t+d$. For each step $i\in[1,d]$, instead of sampling from $\text{SoftMax}(\mathbf{W}\mathbf{o}_{t+i})$, we resample from $\text{SoftMax}(\mathbf{W}\mathbf{o}_{t+i}-\text{n}_{2}({F_{i}>0}))$ where $\text{n}_{2}({F_{i}>0})$ is a vector where each component $j$ is $\text{n}_{2}$ if $F_{i,j}>0$ and 0 otherwise. The addition of $\text{n}_{2}({F_{i}>0})$ disables the possibility of sampling the same token at the same position observed in unsuccessful beams, thus increasing diversity.
It is worth noting that as we sample from the reference LLM and rank responses directly or via the critic, block sampling ensures a small Kullback-Leibler (KL) divergence from the original LLM without explicitly adding a KL regularizer into our objective, preserving coherence and flow; see [14].
6 Experiments
Baselines. We evaluate the helpfulness (task cost) and harmlessness (safety cost) of our method on four models with varying safety alignment levels: the Alpaca-7B model [67], Vicuna-7B [68], Llama3.1-8B-Instruct [69] and the safety-aligned Beaver-v3-7B model [70]. We compare InferenceGuard to the following state-of-the-art test-time alignment methods: Best-of-N (BoN), beam search, and more recent advances ARGS [43] and RE-Control [42] that combines token probabilities with reward scores predicted by pre-trained reward model or value network during the decoding process. These test-time alignment methods were originally designed to maximize rewards without considering safety. To ensure a fair and meaningful comparison, we extend them to also align for safety through the Lagrangian-based approach and the safety augmentation approach. This helps us evaluate our performance against other algorithms and highlights the importance of safety augmentation over Lagrangian approach for effectively balancing rewards and constraint satisfaction.
For beam search and Best-of-N (BON), we select solutions with $c_{\text{task}}+\lambda\mathcal{C}_{\text{safety}}$ where $\lambda$ is the Lagrangian multiplier. Similarly, we extend ARGS so that token selection follows: $-\omega\pi(t|\cdot)+c_{\text{task}}+\lambda\mathcal{C}_{\text{safety}}$, with $\omega$ adjusting the influence of the reference policy. We also considered state augmentation for ARGS and RE-Control but found it ineffective. Since these methods decode token-by-token, they cannot recover once $\text{z}_{t}$ flips the sign, and before that, $\text{z}_{t}$ has no influence. Thus, we excluded it from our evaluation. To further strengthen our comparison, we introduce safety-augmented versions of BoN and beam search as additional baselines.
Datasets. We evaluate across three widely recognized safety assessment benchmarks with varying sensitivity: 1) PKU-SafeRLHF includes 37,400 training samples and 3,400 testing samples of safety-critical instructions for reward and cost alignment; 2) HEx-PHI consists of 330 harmful instructions across 11 safety-relevant categories; and 3) HH-RLHF contains 112,000 training samples and 12,500 testing samples of high-safety prompts with strong human preferences. To train the critic network, we construct a dataset by generating five responses per training set prompt from the base model. For HEx-PHI evaluation, we use a value network trained on HH-RLHF, due to the limited size of HEx-PHI data.
Evaluation Metrics. We assess the performance using several metrics: the Average Reward is computed using the reward models from [43] for Llama2-based models and [71] for the Llama3-based model as $-c_{\text{task}}$ on the complete response to reflect helpfulness, where a higher reward indicates better helpfulness; the Average Cost is evaluated with the cost model from [28] and [71], for Llama2 and Llama3 models respectively, as $\mathcal{C}_{\text{safety}}$, indicating harmfulness, with higher cost values reflecting more harmful outputs; the Safety Rate is the proportion of responses where the cumulative cost does not exceed the safety budget $z_{t=0}=10$, and is given by Safety $\text{Safety Rate}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}(\mathcal{C}^{i}_{\text{test}}\leq z_{t=0})$, where $N$ is the number of prompts; and Inference Time refers to the inference time taken to generate one complete response in seconds.
Results. We present our main results in Figure 2 and additional ones in Tables 2, 3 and 4 in Appendix D. InferenceGuard achieves the highest safety rates with all models (reaching up to 94.46% on Alpaca, 98.45% on Llama-3.1-8B-Instruct, 98.97% on Vicuna, and 100% on Beaver), with minimal latency overhead in comparison to baselines (see also Table. 9). With Beaver, our method dominates the Patero front, achieving the highest rewards without any unsafe responses. Although Lagrangian methods can have a reasonable average cost, they fail to satisfy the safety constraints. Moreover, they are too safe on already safe answers, hindering their rewards. The RE-Control intervention method underperforms for all models in our setting. ARGS can provide safe answers but with very poor rewards because most answers are very short to avoid breaking the safety constraint. Among augmented safety methods, Best-of-N’s inability to leverage intermediate signals and beam search’s blindness to past mistakes lead to more unsafe answers and inferior performance compared to InferenceGuard.
Figure 3 provides a better view of the reward and safety distributions. The figure shows that InferenceGuard consistently achieves higher rewards while maintaining low cumulative costs, outperforming other methods. Its cumulative cost stays just under the safety budget while maximizing the reward as suggested by our theoretical contributions. Finally, we observe that the trained critic helps to better guide the search on intermediate trajectories in both the unaligned Alpaca-7B model and the safety-aligned Beaver-7B model. We further provide ablation studies, limitations, latency overhead, critic training details, and qualitative comparisons of generated answers in Appendix D.
Figure 3: Reward and cost distributions of responses generated from Alpaca-7B (top) and Beaver-v3 (bottom) on PKU-SafeRLHF data. The left y-axis indicates the reward, while the right y-axis shows the cumulative cost. InferenceGuard outperforms baselines, both in terms of rewards and safety costs.
7 Conclusion
We introduced InferenceGuard, a novel inference-time alignment method that aims to ensure LLMs generate safe responses almost surely w.r.t. given safety cost model. We extended prior safety-augmented MDP theorems into the latent space of LLMs and conducted a new analysis. Our results demonstrated that InferenceGuard significantly outperforms existing test-time alignment methods, achieving state-of-the-art safety versus reward tradeoff results.
Limitations. Our almost sure safety guarantees rely on the quality of the given safety cost model and, hence, improving its safety assessment ability remains a subject for future investigation. We also plan to improve the algorithm’s efficiency further and generalize our setting to cover jailbreaking. While our method is, in principle, extendable to jailbreaking settings, we aim to analyze whether our theoretical guarantees still hold.
Appendix A Additional Related Work
Safe RL: Safe RL employs the cMDP framework [72] to enforce safety constraints during exploration and policy optimization. When no prior knowledge is available, methods focus on safe exploration [73, 74, 75, 76, 77]. With prior knowledge, such as environmental data or an initial safe policy, methods learn safe policies using control techniques like Lyapunov stability [78, 79, 80, 81] and reachability analysis [82, 83, 84, 85]. Safety constraints are enforced via Lagrangian or constrained optimization methods [65, 86, 87, 88, 89, 70], but can often lead to suboptimal safety-reward trade-offs. In contrast, our approach extends safety state augmentation [16] to LLMs and latent MDPs to ensure almost sure inference time safety without relying on Lagrangian multipliers.
LLM alignment and safety: Methods for aligning pre-trained LLMs with task-specific data include prompting, guided decoding, and fine-tuning. Among fine-tuning methods, RL from Human Feedback (RLHF) has proven effective, where LLMs are fine-tuned with a learned reward model [2, 17, 11] or directly optimized from human preferences [21, 22, 23, 24, 25, 26, 90]. Recent works have explored fine-tuning for helpful and harmless responses [27, 8], while [28] introduced a safe RL approach incorporating safety cost functions via Lagrangian optimization, requiring model weight fine-tuning. Other safety-focused methods, including machine unlearning [29], DPO with expanded safety zone [91], dual objective DPO with targeted unlearning [92], safety pre-aligned multi-modal LLMs [30], safety-aware model merging,editing [31, 93, 94], prompting-based safety methodologies [32, 95, 96], test-time controllable safety alignment [33], defenses against adversarial attacks and jailbreaking [34, 97, 35, 98, 99, 100, 101, 102, 103], identifying safety critical regions in LLMs [36], safety preserved LoRA fine-tuning [37], alignment using correctional residuals between preferred and dispreferred answers using a small model [104], layer-specific editing of LLMs to ensure safety [105], training a small model to correct the outputs of a large model [106], and identifying safety directions in embedding space [107]. Those methods are either orthogonal, handle a different problem to ours, or can not ensure almost sure safety during inference.
Inference time alignment: The closest literature to ours is inference-time alignment. Those methods offer flexible alternatives to fine-tuning LLMs, as they avoid modifying the model weights. A common approach is guided decoding, which steers token generation based on a reward model. In particular, [43, 44, 45] perform this guided decoding through scores from the reward model whereas Han et al. [46], Mudgal et al. [14], Kong et al. [42] use a value function that is trained on the given reward model. These inference-time alignment methods build on previous works like [40, 108, 109, 110, 111, 112], which guide or constrain LLMs towards specific objectives. Other safety-focused inference-time methods include, reverse prompt contrastive decoding [47], adjusting model hidden states combined with guided decoding [113, 114], cross-model guidance through safety steering vectors [115], adjusting logits based on self-evaluation [116, 35], soft prompt-tuned detoxifier based decoding [49], jailbreak-targeted inference-time interventions[53, 117], speculating decoding using safety classifier [50], comparing the cosine similarity between the target prompt and a set of harmful and benign prompts to classify for refusal [118], decoding using two fine-tuned models (one trained for safety and other as an adversary) [119, 120], and opposite prompt-based contrastive decoding [48]. Compared to those methods, we are the first to achieve almost sure safe alignment with strong empirical results. Operating in the latent space enables us to train smaller, inference-efficient critics while optimally balancing rewards and safety constraints (see Section 6) without introducing additional parameters like Lagrangian multipliers.
Appendix B Theoretical Analysis
For our theoretical results, we consider a similar setup to that of Sootla et al. [16], Hernández-Lerma and Muñoz de Ozak [15] but with a discrete action space. Consider an MDP $M=\{S,A,P,c,\gamma_{c}\}$ with a discrete, non-empty, and finite action set for each $s$ defined as $A(s)$. The set
$K=\{(s,a)\mid s\in S,a\in A(s)\}$ (10)
defines the admissible state-action pairs and is assumed to be a Borel subset of $S\times A$. A function $u$ is inf-compact on $K$ if the set $\{a\in A(s)\mid u(s,a)\leq r\}$ is compact for every $s\in S$ and $r\in\mathbb{R}$. Note that, since the action space is finite and discrete every function $u$ is inf-compact on $K$. A function $u$ is lower semi-continuous (l.s.c.) in $S$ if for every $s_{0}\in S$ we have
$\liminf_{s\to s_{0}}u(s)\geq u(s_{0}).$
Let $L(S)$ denote the class of all functions on $S$ that are l.s.c. and bounded from below.
For a given action, $a\in A(s)$, a distribution $P(y\mid s,a)$ is called weakly continuous w.r.t. $s$, if for any function $u(s)$, continuous and bounded w.r.t. $s$ on $S$, the map
$(s,a)\mapsto\int_{S}u(y)P(dy\mid s,a)$
is continuous on $K$ for a given $a$.
We also make the following assumptions:
B1. The function $c(s,a)$ is bounded, measurable on $K$, nonnegative, lower semi-continuous w.r.t. $s$ for a given $a\in A(s)$;
B2. The transition law $P$ is weakly continuous w.r.t. $s$ for a given $a\in A(s)$;
B3. The set-value function map $s:A(s)$ satisfies the following, $\forall s_{0}\in\mathcal{S}$, there exists a $\epsilon>0$, such that $\forall x$ satisfying $\|x-x_{0}\|\leq\epsilon$, $A(x)=A(x_{0})$
Note that, the assumptions B1-B3, share a similar essence to that of the Assumptions 2.1-2.3 in Hernández-Lerma and Muñoz de Ozak [15] and B1-B3 in Sootla et al. [16] but suited for a discrete action space. In particular, Assumption B3, is similar to the lower semi continuity assumption on the set-value function map $A(s)$ taken in Sootla et al. [16], Hernández-Lerma and Muñoz de Ozak [15] but modified for a discrete action space.
Our first goal is to recreate Hernández-Lerma and Muñoz de Ozak [15, Lemma 2.7] for our discrete action setting. Let $\Pi$ denote the set of functions from $S\to A$.
Lemma 1.
  • (a) If Assumption B3 holds and $v(s,a)$ is l.s.c. w.r.t. $s$ for any given $a\in A(s)$ and bounded from below on the set $K$ (see Equation 10), then the function
    $v^{*}(s):=\inf_{a\in A(s)}v(s,a)$
    belongs to $L(S)$ and, furthermore, there is a function $\pi\in\Pi$ such that
    $v^{*}(s)=v(s,\pi(s))\quad\forall s\in S.$
  • (b) If the Assumptions B1-B3 hold, and $u\in L(S)$ is nonnegative, then the (nonnegative) function
    $u^{*}(s):=\inf_{a\in A}\left[c(s,a)+\int_{S}u(y)P(dy\mid s,a)\right]$
    belongs to $L(S)$, and there exists $\pi\in\Pi$ such that
    $u^{*}(s)=c(s,\pi(s))+\int_{S}u(y)P(dy\mid s,\pi(s))\quad\forall s\in S.$
  • (c) For each $n=0,1,\dots$, let $v_{n}$ be a l.s.c. function, bounded from below. If $v_{n}\to v_{0}$ as $n\to\infty$, then
    $\lim_{n\to\infty}\inf_{a\in A(s)}v_{n}(s,a)=\inf_{a\in A(s)}v_{0}(s,a)\quad\forall s\in S.$
Proof.
For part a, note that, we have $v(s,a)$ is l.s.c. w.r.t. $s$ for any given $a$. This implies from the definition of lower semi-continuity for any $s_{0}\in S$ and $a\in A(s_{0})$, if $v(s_{0},a)>y$, then there exists a $\epsilon>0$, s.t. $\forall s$ satisfying $\|s-s_{0}\|\leq\epsilon$, $v(s,a)>y$.
Assume for some $s_{0}\in S$ and $y$, the function $\inf_{a\in A(s_{0})}v(s_{0},a)$ satisfies,
$\inf_{a\in A(s_{0})}v(s_{0},a)>y\quad\Rightarrow v(s_{0},a)>y\quad\forall a\in A(s_{0})$ (11)
Using Assumption B3, we have $A(s)=A(s_{0})$ for $\|s-s_{0}\|\leq\epsilon$. Moreover, using the fact that $v(s,a)$ is l.s.c. at a given $a$, we have if $v(s_{0},a)>y$, $\forall a\in A(s_{0})$ we have,
$v(s,a)>y\quad\forall\|s-s_{0}\|\leq\epsilon,\forall a\in A(s)$ (12)
Since, this holds for all $a\in A(s)$, it also holds for
$\inf_{a\in A(s)}v(s,a)>y\quad\forall\|s-s_{0}\|\leq\epsilon$ (13)
This proves the lower semi continuity of $v^{*}(s)=\inf_{a\in A(s)}v(s,a)$. Further, due to the discrete nature of $A(s)$, the $\inf_{a\in A(s)}v(s,a)$ is always attained by an action $\pi(s)\in A(s)$. Hence, there exists a function $\Pi:S->A(s)$, such that,
$v^{*}(s)=\inf_{a\in A(s)}v(s,a)=v(s,\pi(s))\quad\forall s\in S.$ (14)
For part b), note that $c(s,a)+\int u(y)P(dy\mid s,a)$ is l.s.c. w.r.t. $s$ for an given $a$, based on Assumptions B1-B2. Hence, using part a) we have $\forall s\in S$,
$u^{*}(s):=\inf_{a\in A}\left[c(s,a)+\int_{S}u(y)P(dy\mid s,a)\right]=c(s,\pi(s))+\int_{S}u(y)P(dy\mid s,\pi(s))\in L(S)$ (15)
for some $f\in\Pi$.
For part c), we begin by defining $l(s)=lim_{n\to\infty}\inf_{a\in A(s)}v_{n}(s,a)$. Note that, since $\{v_{n}\}$ is an increasing sequence, we have for any $n$
$\inf_{a\in A(s)}v_{n}(s,a)\leq\inf_{a\in A(s)}v_{0}(s,a)$ (16)
This implies,
$l(s)\leq\inf_{a\in A(s)}v_{0}(s,a)=v_{0}^{*}(s)$ (17)
Next, we define for any $s\in S$,
$A_{n}:=\{a\in A(s)|v_{n}(s,a)\leq v_{0}^{*}(s)\}$ (18)
We note that $A_{n}$ are compact sets as $A$ is finite and discrete. Further, note that $A_{n}$ is a decreasing sequence converging to $A_{0}$ (compact, decreasing and bounded from below by $A_{0}$). Also, note that
$A_{1}\supset A_{2}\supset A_{3}\cdots\supset A_{0}$ (19)
We consider the sequence $\{a_{n}\}$ where $a_{n}\in A_{n}$ and $a_{n}$ satisfies,
$v_{n}(s,a_{n})=\inf_{a\in A(s)}v_{n}(s,a)\leq\inf_{a\in A(s)}v_{0}(s,a)\leq v_{0}^{*}(s)$ (20)
This sequence $\{a_{n}\}$ belongs to the compact space $\cup_{n=1}^{\infty}A_{n}=A_{1}$, hence it has convergent subsequence $\{a_{n_{i}}\}$ converging to $\cup_{n=1}^{\infty}A_{n}=A_{1}$.
$a_{n_{i}}$ $\in A_{n_{i}}=\cap_{n\leq n_{i}}A_{n}$ (21)
$a_{0}$ $\in\cap_{n\leq\infty}A_{n}=A_{0}$ (22)
Since, the converging sequence $a_{n_{i}}\to a_{0}$ belongs to the discrete, compact space, there exisits a $N_{i}$, such that for all $n_{i}\geq N_{i}$, $a_{n_{i}}=a_{0}$. Further, using the increasing nature of $v_{n}$, we have,
$v_{n_{i}}(s,a_{n_{i}})\geq v_{n}(s,a_{n_{i}})\quad\forall n_{i}\geq n$ (23)
As $i\to\infty$, this implies,
$\lim_{i\to\infty}v_{n_{i}}(s,a_{n_{i}})\geq v_{n}(s,a_{0})$ (24)
$\lim_{i\to\infty}\inf_{a\in A(s)}v_{n_{i}}(s,a)\geq v_{n}(s,a_{0})$ (25)
$l(s)\geq v_{n}(s,a_{0})$ (26)
As $v_{n}\to v_{0}$, $l(s)\geq v_{0}(s,a_{0})=v_{0}^{*}(s)$.
B.1 Optimality Equation
In this section, we characterize the solution to the Bellman (Optimality) equation. We begin by recalling the Bellman operator:
$Tv(s)=\min_{a\in A(s)}\left\{c(s,a)+\gamma\int v(y)P(dy\mid s,a)\right\}.$
To state our next result we introduce some notation: Let $L(S)^{+}$ be the class of nonnegative and l.s.c. functions on $S$, and for each $u\in L(S)^{+}$ by Lemma 1(b), the operator $T$ maps $L(S)^{+}$ into itself. We also consider the sequence $\{v_{n}\}$ of value iteration (VI) functions defined recursively by
$v_{0}(S):=0,\quad\text{and}\quad v_{h}:=Tv_{h-1}\quad\text{for}\quad h=1,2,\dots$
That is, for $h\geq 1$ and $s\in S$,
$v_{h}(s):=\min_{a\in\mathcal{A}(s)}\left(c(s,a)+a\int_{S}v_{h-1}(y)P(dy\mid s,a)\right).$
Note that, by induction and Lemma 1(b) again, $v_{h}\in L(S)^{+}$ for all $h\geq 0$. From elementary Dynamic Programming [121, 122, 123], $v_{h}(s)$ is the optimal cost function for an $h$-stage problem (with "terminal cost" $v_{0}(s)=0$) given $s_{0}=s$; i.e.,
$v_{h}(s)=\inf_{\pi}V_{h}(\pi,s),$
where, $\Pi$ is the set of policies and $V_{H}(\pi,s)$ denotes the value function for the $H-$stage problem:
$V_{H}(\pi,s_{0})=\mathbb{E}_{\pi}\left[\sum_{h=0}^{H-1}\gamma^{h}c(s_{h},a_{h})\right].$
Here, $\mathbb{E}_{\pi}$ stands for the expectation with actions sampled according to the policy $\pi$ and the transitions $P$. For $H\to\infty$, let the value functions be denoted as follows:
$V(\pi,s_{0})=\mathbb{E}_{\pi}\left[\sum_{h=0}^{\infty}\gamma^{h}c(s_{h},a_{h})\right],$
and
$V^{*}(s)=\inf_{\pi}V(\pi,s).$
We want to prove similar results to that of Hernández-Lerma and Muñoz de Ozak [15, Theorem 4.2] on the optimality of the Bellman operator, however in the discrete action setting. In particular, we want to show the following theorem
Theorem 2.
Suppose that Assumptions B1-B3 hold, then:
  • (a)
    $v_{h}\to V^{*}$; hence
  • (b)
    $V^{*}$ is the minimal pointwise function in $L(S)^{+}$ that satisfies
    $V^{*}=TV^{*}$
Proof.
We follow a similar proof strategy to that of Hernández-Lerma and Muñoz de Ozak [15, Theorem 4.2].
To begin, note that the operator $T$ is monotone on $L(S)^{+}$, i.e., $u>v$ implies $Tu>Tv$. Hence $\{v_{h}\}$ forms a nondecreasing sequence in $L(S)^{+}$ and, therefore, there exists a function $u\in L(S)^{+}$ such that $v_{h}\to u$. This implies (by the Monotone Convergence Theorem) that
$c(s,a)+a\int v_{h-1}(y)P(dy\mid s,a)\to c(s,a)+a\int u(y)P(dy\mid s,a),$
Using Lemma 1(c), and $v_{h}=\inf_{a\in A(s)}\{c(s,a)+a\int v_{h-1}(y)P(dy\mid s,a)\}$ yields
$\lim_{h\to\infty}\inf_{a\in A(s)}\{c(s,a)+a\int v_{h-1}(y)P(dy\mid s,a)\}$ $=\inf_{a\in A(s)}\{c(s,a)+a\int u(y)P(dy\mid s,a)\},$
$\lim_{h\to\infty}v_{h}$ $=Tu,$
$u$ $=Tu.$
This shows $v_{h}\to u$, such that $u\in L(S)^{+}$ satisfies the Optimality equation.
Next, we want to show $u=V^{*}$. Using that $u\geq Tu$, and by Lemma 1(b), we have that there exists $\pi\in\Pi$, a stationary policy that satisfies
$u(s)$ $\geq\inf_{a\in A(s)}\{c(s,a)+a\int u(y)P(dy\mid s,a)\}\quad\forall s$ (27)
$\geq c(x,\pi)+\alpha\int u(y)P(dy\mid s,\pi)\quad\forall s.$ (28)
Applying the $T$ operator iteratively, we have
$u(s)$ $\geq T^{H}u(s)\quad\forall s,H$ (29)
$\geq\mathbb{E}_{\{s_{h}\},\pi}[\sum_{h=0}^{H-1}\alpha^{h}c(s_{h},\pi)]+\alpha^{H}\int u(y)P^{H}(dy\mid s,\pi)\quad\forall s,H,$ (30)
where $P^{H}(B\mid s,\pi)=P(\{s_{H}\in B\})$ denotes the $H$-step transition probability of the Markov chain $\{s_{h}\}$ (see Hernández-Lerma and Muñoz de Ozak [15, Remarks 3.1,3.2]). Therefore, since $u$ is nonnegative,
$u(s)$ $\geq\mathbb{E}_{\{s_{h}\},\pi}[\sum_{h=0}^{H-1}\alpha^{h}c(s_{h},\pi)]\quad\forall s,H,$ (31)
Letting $H\to\infty$, we obtain
$u(s)\geq V(\pi,s)\geq V^{*}(s)\quad\forall s.$
Next, note that,
$v_{h}(s)$ $=\inf_{\pi\in\Pi}V_{h}(\pi,s)\leq V_{h}(\pi,s)\quad\forall s,h,\pi$ (32)
and letting $h\to\infty$, we get
$u(s)\leq V(\pi,s)\quad\forall s,\pi.$
This implies $u(s)\leq V^{*}(s)$. We have thus shown that $u=V^{*}$. Further, if there is another solution $u^{\prime}$ satisfying $u^{\prime}=Tu^{\prime}$, it holds that $u^{\prime}\geq V^{*}$, Hence, $V^{*}$ is the minimal solution. ∎
B.2 Limit of a sequence of MDPs
Consider a sequence of Markov Decision Processes (MDPs) $M_{n}=\{S,A,P,c_{n},\gamma_{n}\}$, where, without loss of generality, we write $c_{n},c_{\infty}$ and $M_{n},M_{\infty}$ with corresponding value functions $\{V_{n}^{*}\}_{n=0}^{\infty}$:
$V_{n}(\pi,s_{0})=\mathbb{E}^{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}c_{n}(s_{t},a_{t})\right],\quad V_{n}^{*}(s)=\inf_{\pi}V_{n}(\pi,s).$
The "limit" value functions (with $n=\infty$) are still denoted as follows:
$V(\pi,s_{0})=\mathbb{E}^{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}c(s_{t},a_{t})\right],\quad V^{*}(s)=\inf_{\pi}V(\pi,s).$
We also define the sequence of Bellman operators
$T_{n}v(s)=\min_{a\in A(s)}\left\{c_{n}(s,a)+\gamma\int v(y)P(dy\mid s,a)\right\},$
$Tv(s)=\min_{a\in A(s)}\left\{c(s,a)+\gamma\int v(y)P(dy\mid s,a)\right\}.$
In addition to the previous assumptions, we make an additional one, and modify Assumption B1:
B1’ For each $n$, the functions $c_{n}(s,a)$ are bounded, measurable on $\mathcal{K}$, nonnegative, lower semi-continuous w.r.t. $s$ for a given $a\in A(s)$;
B4 The sequence of cost functions $\{c_{n}(s,a)\}_{n=0}^{\infty}$ converges to $c$, i.e, $c_{n}\uparrow c$.
For each $n$, the optimal cost function $V^{*}(s)$ is the bounded function in $L(S)^{+}$ that satisfies the Optimality equation in 2 :
$V_{n}^{*}=T_{n}V_{n}^{*},$
Theorem 3.
The sequence $V_{n}^{*}$ is monotone increasing and converges to $V^{*}$.
Proof.
To begin with, note that since $c_{n}\uparrow c$, it is clear that $V_{n}^{*}$ is an increasing sequence in $L(S)^{+}$, and therefore, there exists a function $u\in L(S)^{+}$ such that $V_{n}^{*}\to u$.
Moreover, from Lemma 1(c), letting $n\to\infty$, we see that $u=Tu$, i.e., $u$ satisfies the optimality equation. This implies that $u\geq V^{*}$, since, by Theorem 2, $V^{*}$ is the minimal solution in $L(X)^{+}$ to the optimality equation.
On the other hand, it is clear that $V_{n}^{*}\leq V^{*}$ for all $n$, so that $u\leq V^{*}$. Thus $u=V^{*}$, i.e., $U^{*}=V^{*}$.
B.3 Latent MDP Analysis
[…]
Proof.
We begin by comparing our assumptions to that of the assumptions B1-B4, closely aligned to those used in Hernández-Lerma and Muñoz de Ozak [15], Sootla et al. [16].
To prove a),b) of Definition 5.1 we need to verify that the latent MDP satisfying Assumptions A1-A2 also satisfies Assumptions B1’, B2-B4. According to Assumption A1, we consider bounded costs $\bar{\mathcal{C}}^{n}_{\text{task}}$ continuous w.r.t. state $(\mathbf{h},\mathbf{o},\text{z})$ for a given $y$ with discrete and finite action space ${\mathcal{V}}$, hence Assumptions B1’, B3, and B4 are satisfied. Assumptions B2 and A2 are identical. This proves a), b).
For c), note that the state value function $V(\cdot)$ and latent space value function $\bar{V}(\cdot)$ w.r.t. policy $\bar{\pi}:\bar{{\mathcal{S}}}\rightarrow{\mathcal{V}}$ that acts on the latent space directly and on the original space as $\bar{\pi}(\phi(\cdot)):{\mathcal{S}}\rightarrow{\mathcal{V}}$ are related as follows:
$V(\bar{\pi}(\phi(\cdot)),\mathbf{s}_{0},\text{z}_{0})$ $=\mathbb{E}_{\begin{subarray}{c}\mathbf{s}_{t+1},\text{z}_{t+1}\sim{\mathcal{P}}(\cdot|\mathbf{s}_{t},\text{z}_{t},\text{y}_{t})\\ \text{y}_{t}\sim\bar{\pi}(\cdot|\phi(\mathbf{s}_{t}),\text{z}_{t})\end{subarray}}\Big{[}\sum_{t=0}^{\infty}\gamma^{t}\tilde{\mathcal{C}}^{n}_{\text{task}}(\mathbf{s}_{t},\text{z}_{t},\text{y}_{t})\Big{]}$ (33)
$=\mathbb{E}_{\begin{subarray}{c}\mathbf{s}_{t+1},\text{z}_{t+1}\sim{\mathcal{P}}(\cdot|\mathbf{s}_{t},\text{z}_{t},\text{y}_{t})\\ \text{y}_{t}\sim\bar{\pi}(\cdot|\phi(\mathbf{s}_{t}),\text{z}_{t})\end{subarray}}\Big{[}\sum_{t=0}^{\infty}\gamma^{t}\bar{\mathcal{C}}^{n}_{\text{task}}(\phi(\mathbf{s}_{t}),\text{z}_{t},\text{y}_{t})\Big{]}$ (34)
$=\mathbb{E}_{\begin{subarray}{c}\phi(\mathbf{s}_{t+1}),\text{z}_{t+1}\sim\bar{{\mathcal{P}}}(\cdot|\phi(\mathbf{s}_{t}),\text{z}_{t})\big{)}\\ \text{y}_{t}\sim\bar{\pi}(\cdot|\phi(\mathbf{s}_{t}),\mathbf{o}_{t})\end{subarray}}\Big{[}\sum_{t=0}^{\infty}\gamma^{t}\bar{\mathcal{C}}^{n}_{\text{task}}(\phi(\mathbf{s}_{t}),\text{z}_{t},\text{y}_{t})\Big{]}\Big{|}$ (35)
$=\mathbb{E}_{\begin{subarray}{c}\mathbf{h}_{t+1},\mathbf{o}_{t+1},\text{z}_{t+1}\sim\bar{{\mathcal{P}}}(\cdot|\mathbf{h}_{t},\mathbf{o}_{t},\text{z}_{t})\big{)}\\ \mathbf{h}_{t+1},\mathbf{o}_{t+1}=\phi(\mathbf{s}_{t+1})\\ \text{y}_{t}\sim\bar{\pi}(\cdot|\mathbf{h}_{t},\mathbf{o}_{t})\end{subarray}}\Big{[}\sum_{t=0}^{\infty}\gamma^{t}\bar{\mathcal{C}}^{n}_{\text{task}}(\mathbf{h}_{t},\mathbf{o}_{t},\text{z}_{t},\text{y}_{t})\Big{]}\Big{|}_{\mathbf{h}_{0},\mathbf{o}_{0}=\phi(s_{0})}$ (36)
$=\bar{V}(\bar{\pi},\mathbf{h}_{0},\mathbf{o}_{0},z_{0})\Big{|}_{\mathbf{h}_{0},\mathbf{o}_{0}=\phi(s_{0})}$ (37)
Hence, we can show $\bar{\pi}_{n}^{*}$ is optimal for $V_{n}(\cdot)$ as follows:
$V_{n}(\bar{\pi}_{n}^{*},\mathbf{s},\text{z})$ $=\bar{V}_{n}(\bar{\pi}_{n}^{*},\mathbf{h},\mathbf{o},\text{z})\Big{|}_{\mathbf{h},\mathbf{o}=\phi(s)}$ (38)
$=\min_{\bar{\pi}}\bar{V}_{n}(\bar{\pi},\mathbf{h},\mathbf{o},\text{z})\Big{|}_{\mathbf{h},\mathbf{o}=\phi(s)}$ (39)
$=\min_{\bar{\pi}}V_{n}(\bar{\pi}(\phi(\cdot)),\mathbf{s},\text{z})$ (40)
$=\min_{\pi}V_{n}(\pi,\mathbf{s},\text{z})$ (41)
Here, the minimization of $\pi$ is over set of all policies covered by $\bar{\pi}(\phi(\cdot))$ and we show that $\bar{\pi}_{n}^{*}(\phi(\cdot))$ is the optimal policy for the original space over this set of policies.
[…]
Proof.
We first note that if any trajectory with infinite cost has a finite probability, the cost would be infinite. Hence, all the trajectories with finite/positive probability have finite costs. This implies, the finite cost attained by $\pi^{\star}$ w.r.t. Equation 7 implies the satisfaction of constraints (Equation 9) almost surely (i.e. with probability 1). Combined with the fact that the policy $\pi^{\star}$ was obtained by minimizing the exact task cost as in Equation 9, Definition 5.1 follows. ∎
Appendix C Algorithmic details
Input: Initial prompt $s_{0}$, beam depth $d$, $N$ number of beams, max depth $D$, top $K$ beams, $M$ max retry, reference policy $\pi_{\text{ref}}$
Initialize the beam $B\leftarrow\{s_{0}\}$
for $\frac{D}{d}$ iterations do
      $\forall i\in\{1,\dots d\},j\in\{1,...,|\mathcal{V}|\},F_{i,j}\leftarrow$ 0, $\pi_{\text{pen}}\leftarrow\pi_{\text{ref}}$
       for $M$ rounds do
            $B_{new}\leftarrow$ Generate $N$ continuations of length $d$ from $B$ with $\pi_{\text{pen}}$ and $F$
[…]
             if $\exists i,E_{i}>0$ or last round then
                  $B\leftarrow B_{\text{new}}$ ;
                   break
             end if
            $F\leftarrow F+token\_frequency(B_{\text{new}})$
             $\pi_{\text{pen}}\leftarrow\pi_{\text{pen}}/\text{n}_{2}({F>0})$
       end for
      Keep top $K$ beams in $B$ according to their scores E
end for
return Best trajectory in $B$ according to E
Algorithm 1 InferenceGuard
Here, $\pi_{\text{pen}}$ is the penalized policy that for each step $t+i\in[t,t+d]$, samples from $\text{SoftMax}(\mathbf{W}\mathbf{o}_{t+i}-\text{n}_{2}({F_{i}>0}))$ where $\text{n}_{2}({F_{i}>0})$ is a vector where each component $j$ is $\text{n}_{2}$ if $F_{j,i}>0$ and 0 otherwise. This avoids sampling the same token at the same position observed in unsuccessful beams, thus increasing diversity.
If the cost functions allow intermediate evaluations, we evaluate a beam $y_{t},\cdots,y_{t+d}$ but using our augmented cost function:
$\text{E}_{\text{inter}}(y_{t},\cdots,y_{t+d})=\begin{cases}\gamma^{T}\bar{c}_{\text{task}}({\bm{h}}_{t+d},{\bm{o}}_{t+d})&\text{z}_{t+d}>0\\ n&\text{otherwise}.\end{cases}$
When we only have a critic, we use:
$\text{E}_{\text{critic}}(y_{t},\cdots,y_{t+d})=$$\begin{cases}\gamma^{T}\bar{c}_{\text{task}}({\bm{h}}_{t+d})&t+d=T\text{ and }\text{z}_{t+d}>0\\ n&t+d=T\text{ and }\text{z}_{t+d}\leq 0\\ f^{2}_{\bm{\theta}}({\bm{h}}_{t+d},{\bm{o}}_{t+d},\text{z}_{t+d})&f^{1}_{\bm{\theta}}({\bm{h}}_{t+d},{\bm{o}}_{t+d},\text{z}_{t+d})>0.5\\ n&\text{otherwise}.\end{cases}$
If we can do both, as $\text{z}_{t}$ only decreases overtime, the critic head predicting the safety would only act as an additional filter. We introduce another hyper-parameter $\eta$ to balance the confidence in the second head of the critic predicting the future cost:
$\text{E}_{\text{mix}}(y_{t},\cdots,y_{t+d})=\begin{cases}\gamma^{T}\bar{c}_{\text{task}}(\mathbf{h}_{t+d},\bm{o}_{t+d})&t+d=T\text{ and }\text{z}_{t+d}>0\\ n&t+d=T\text{ and }\text{z}_{t+d}\leq 0\\ \gamma^{T}\bar{c}_{\text{task}}({\bm{h}}_{t+d},\bm{o}_{t+d})+\eta f^{2}_{\bm{\theta}}({\bm{h}}_{t+d},{\bm{o}}_{t+d},\text{z}_{t+d})&f^{1}_{\bm{\theta}}({\bm{h}}_{t+d},{\bm{o}}_{t+d},\text{z}_{t+d})>0.5\text{ and }\text{z}_{t+d}>0\\ n&\text{otherwise}.\end{cases}$
Appendix D Experimental Details
D.1 Experiment Setup
All the experiments are conducted on our internal cluster using Python 3.11 and bfloat16 mixed-precision training. The setup used a single high-memory compute device, and the inference process can be reproduced on any similar hardware with at least 64 GB of on-device memory available. We assess safe inference across three datasets that vary in their safety sensitivity: 1) PKU-SafeRLHF with safety-critical instructions for reward and cost alignment, 2) HEx-PHI: mixed-sensitivity health and philosophy instructions with challenging edge cases, and 3) HH-RLHF with high-safety prompts with strong human preferences. Our experiments involved four major language models with varying safety alignment levels: the unaligned Alpaca-7B model [^5] https://huggingface.co/PKU-Alignment/alpaca-7b-reproduced[67], the safety-aligned Beaver-v3-7B model[70] [^6] https://huggingface.co/PKU-Alignment/beaver-7b-v3.0, the unaligned Vicuna-7B[68] [^7] https://huggingface.co/lmsys/vicuna-7b-v1.5 and Llama3.1-8B-Instruct models [69] [^8] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct. We employed two sets of reward models: 1) Llama2-based reward model llama-7b-rm[43] [^9] https://huggingface.co/argsearch/llama-7b-rm-float32 and cost model beaver-7b-unified-cost [70] [^10] https://huggingface.co/PKU-Alignment/beaver-7b-unified-cost, and 2) Llama-3-based reward model and cost model QRM-LLama-3.1-8B[71] [^11] https://huggingface.co/nicolinho/QRM-Llama3.1-8B during the training stage for critic network, test-time inference and evaluation stages.
D.2 Experiment Settings and Results
The baseline methods we compare against include BoN, Beam Search, RE-Control[42], and ARGS[43]. For BoN, we use a strategy where the top N outputs are sampled, and the best result is selected based on a predefined criterion. Similarly, Beam Search explores multiple beams during the search process and selects the best output based on a beam-width parameter. In RE-Control, an MLP network is employed as the value network to intervene in the decision-making process, guiding the generation through reinforcement learning[42]. ARGS, on the other hand, implements a logits-based greedy token-wise search strategy, where tokens are generated sequentially based on the maximum likelihood of the next token[43].
Given the limited research on safe alignment with inference-time methods, we adapt these baseline methods to enable safe inference using the Lagrangian and safety augmentation approaches, ensuring a fair comparison. To this end, we incorporate a Lagrangian multiplier term or a safety augmentation based on their open-source implementations to enable safety-aligned inference. Notably, BoN and Beam Search utilize a form of blocking sampling, while ARGS and RE-Control employ token sampling methods.
In the Lagrangian-based method setup, we modify the inference process of the baseline methods by incorporating a Lagrangian multiplier, using the following score function: $c_{\text{task}}+\lambda\mathcal{C}_{\text{safety}}$ where $\lambda$ is the Lagrangian multiplier, that controls the influence of the safety cost score $\mathcal{C}_{\text{safety}}$. For the main experiment results, we set $\lambda=5$ uniformly and fix the sampling parameters across all baseline methods and InferenceGuard to ensure a fair comparison. Each method also has additional specific settings, as detailed below:
Method Sample Width $(d)$ Num Samples $(N)$ Other Parameters
ARGS 1 128 N/A
RE-Control 1 N/A n steps = 30, step size = 0.5
BoN 32 64, 128, 256 N/A
Beam Search 32 64, 128, 256 D=128, K=$N/4$
InferenceGuard 32 64, 128, 512 M=2, D=128, K=$N/4$
Table 1: Hyperparameters for Baselines and InferenceGuard.
We compare the performance of baseline methods with lagrangian-based methods and safety augmentation methods, under these hyperparameter settings, and summarize the result in Table 2 for Alpaca-7B and Beaver-7B, Table 3 for Vicuna-7B, and Table 4 for LLaMA-3.1-8B-Instruct. When sampling from the base model, we used a temperature of 1.0 without top_k nor top_p.
Table 2: Performance Comparison using Alpaca-7B and Beaver-7B, evaluated by reward models llama-7b-rm and beaver-7b-unified-cost, on Dataset PKU-SafeRLHF $N=128$, $\lambda=5$
Method Average Reward Safety Rate Inference Time (s)
Alpaca-7B Base 6.15 ($\pm$ 1.51) 29.47% 0.7
RE-Control 6.2 ($\pm$ 1.56) 29.5% 1.13
RE-Control + Lagrangian multiplier 6.19 ($\pm$ 1.50) 29.7% 1.33
Best-of-N + Lagrangian multiplier 5.31 ($\pm$ 1.62) 49.37% 8.0
Best-of-N + Augmented safety 7.22 ($\pm$ 1.90) 58.57% 9.75
Beam search + Lagrangian multiplier 6.58 ($\pm$ 1.95) 50.19% 14.5
Beam search + Augmented safety 8.29 ($\pm$ 2.02) 58.89% 14.6
ARGS $\omega=2.5$ 6.74 ($\pm$ 1.70) 28.19% 58.58
ARGS + Lagrangian multiplier $\omega=2.5$ 3.21 ($\pm$ 1.59) 75.8% 63.42
ARGS + Cost Model $\omega=2.5$ 0.19 ($\pm$ 1.65) 81.6% 55.73
InferenceGuard 7.08 ($\pm$ 2.49) 88.14% 16.25
InferenceGuard with Critic 6.50 ($\pm$ 2.50) 94.46% 15.07
Beaver-7B-v3 Base 5.83 ($\pm$ 1.62) 75.89% 0.8
RE-Control 5.9 ($\pm$ 1.56) 75.9% 1.33
RE-Control + Lagrangian multiplier 5.91 ($\pm$ 1.50) 75.9% 1.73
Best-of-N + Lagrangian multiplier 6.58 ($\pm$ 1.89) 85.7% 14.5
Best-of-N + Augmented safety 9.05 ($\pm$ 1.59) 97.21% 9.62
Beam search + Lagrangian multiplier 8.63 ($\pm$ 1.80) 87.35 % 16.0
Beam search + Augmented safety 10.31 ($\pm$ 1.37) 97.36% 9.75
ARGS $\omega=2.5$ 6.72 ($\pm$ 1.83) 78.5% 67.15
ARGS $\omega=2.5$ + Lagrangian multiplier 2.26 ($\pm$ 1.56) 81% 72.58
ARGS $\omega=2.5$ + Cost Model 0.01 ($\pm$ 1.37) 98.4% 64.28
InferenceGuard 10.26 ($\pm$ 1.42) 99.7% 9.95
InferenceGuard with Ot Critic 10.27 ($\pm$ 1.50) 100% 9.76
Table 3: Performance Comparison using Vicuna-7B, evaluated by reward models llama-7b-rm and beaver-7b-unified-cost, on Datasets HEx-PHI and HH-RLHF $N=128$, $\lambda=5$
Dataset Method Average Reward Safety Rate Inference Time (s)
Vicuna-7B HEx-PHI Base 4.69 ($\pm$ 1.36) 48% 1.2
RE-Control 4.75 ($\pm$ 1.31) 49.33% 1.58
RE-Control + Lagrangian multiplier 4.65 ($\pm$ 1.33) 50.7% 1.75
Best-of-N + Lagrangian multiplier 5.22 ($\pm$ 1.39) 79.3% 9.08
Best-of-N + Augmented safety 6.46 ($\pm$ 1.51) 92.6% 10.04
Beam search + Lagrangian multiplier 5.70 ($\pm$ 1.57) 83% 7.2
Beam search + Augmented safety 7.57 ($\pm$ 1.67) 89.33% 11.59
ARGS $\omega=2.5$ 5.67 ($\pm$ 1.45) 47% 68.23
ARGS $\omega=2.5$ + Lagrangian multiplier 1.72 ($\pm$ 1.96) 93.33% 79.28
ARGS $\omega=2.5$ + Cost Model 0.07 ($\pm$ 1.60) 96% 69.38
InferenceGuard 6.90 ($\pm$ 2.08) 96.67% 11.01
InferenceGuard with Critic 6.99 ($\pm$ 2.1) 96.67% 13.29
Vicuna-7B HH-RLHF Base 5.82 ($\pm$ 1.56) 95% 1.17
RE-Control 5.9 ($\pm$ 1.55) 95.13% 1.45
RE-Control + Lagrangian multiplier 5.85 ($\pm$ 1.50) 95.4% 2.09
Best-of-N + Lagrangian multiplier 6.97 ($\pm$ 2.54) 97.24% 8.32
Best-of-N + Augmented safety 8.33 ($\pm$ 1.95) 98.36% 8.58
Beam search + Lagrangian multiplier 8.05 ($\pm$ 2.25) 97.54% 11.28
Beam search + Augmented safety 9.61 ($\pm$ 2.10) 98.23% 11.84
ARGS $\omega=2.5$ 6.83 ($\pm$ 1.83) 96.2% 70.4
ARGS $\omega=2.5$ + Lagrangian multiplier 2.02 ($\pm$ 1.79 ) 97.54% 73.74
ARGS $\omega=2.5$ + Cost Model 0.46 ($\pm$ 1.73 ) 98.96% 72
InferenceGuard 9.49 ($\pm$ 2.16) 98.97% 11.47
InferenceGuard with Critic 9.48 ($\pm$ 2.16) 98.95% 11.73
Table 4: Performance Comparison using Llama-3.18B-Instruct, evaluated by reward model QRM-Llama-3.1-8B, on Datasets PKU-SafeRLHF and HH-RLHF using $N=128$, $\lambda=5$
Dataset Method Average Reward Safety Rate Inference Time (s)
Llama3.1-8B-Instruct PKU-SafeRLHF Base 0.53 ($\pm$ 0.12) 53.22% 1.13
RE-Control 0.55 ($\pm$ 0.22) 52.7% 1.52
RE-Control + Lagrangian multiplier 0.51 ($\pm$ 0.21) 54.01% 2.11
Best-of-N + Lagrangian multiplier 0.35 ($\pm$ 0.17) 81.29% 5.70
Best-of-N + Augmented safety 0.47 ($\pm$ 0.16) 80.89% 5.71
Beam search + Lagrangian multiplier 0.39 ($\pm$ 0.16) 81.69% 8.60
Beam search + Augmented safety 0.46 ($\pm$ 0.16) 81.43% 8.68
ARGS $\omega=2.5$ 0.55 ($\pm$ 0.21) 52.6% 74.53
ARGS $\omega=2.5$ + Lagrangian multiplier 0.25 ($\pm$ 0.23 ) 73.33% 78.22
ARGS $\omega=2.5$ + Cost Model 0.05 ($\pm$ 0.29) 94.0% 71.61
InferenceGuard 0.44 ($\pm$ 0.13) 90.91% 10.69
Llama3.1-8B-Instruct HH-RLHF Base 0.09 ($\pm$ 0.27) 23.5% 0.82
RE-Control 0.11 ($\pm$ 0.29) 23.3% 1.07
RE-Control + Lagrangian multiplier 0.10 ($\pm$ 0.21) 24.7% 1.26
Best-of-N + Lagrangian multiplier 0.31 ($\pm$ 0.21) 82.37% 5.53
Best-of-N + Augmented safety 0.42 ($\pm$ 0.19) 84.5% 5.61
Beam search + Lagrangian multiplier 0.31 ($\pm$ 0.21) 83.96% 7.12
Beam search + Augmented safety 0.42 ($\pm$ 0.2) 83.15% 7.08
ARGS $\omega=2.5$ 0.52 ($\pm$ 0.29 ) 49.8% 68.9
ARGS $\omega=2.5$ + Lagrangian multiplier 0.24 ($\pm$ 0.19) 87.8% 71.4
ARGS $\omega=2.5$ + Cost Model 0.01 ($\pm$ 0.29 ) 96.01% 64.2
InferenceGuard 0.38 ($\pm$ 0.14) 98.45% 8.11
D.3 Ablation Study
Effect of $\lambda$ and sample size $N$ on baselines
To further evaluate the safety-performance trade-off in baseline methods, we compare decoding strategies across a range of Lagrangian multipliers $\lambda$ and sample sizes $N$ for Best-of-N, Beam Search, and ARGS, under varying $\lambda$ values and sample budgets, as shown in Table 5 for Alpaca-7B and Beaver-7B, and Table 6 for Vicuna-7B. For Beam Search specifically, we apply an alternative technique to enhance sampling diversity by reducing the frequency with which $c_{\text{task}}+\lambda\mathcal{C}_{\text{safety}}$ falls below a threshold; we refer to this variant as Beam Search Lag. + Freq. While safety rates improve as $\lambda$ increases, task reward often degrades sharply, particularly for ARGS. In contrast, InferenceGuard achieves a more favorable balance without requiring fine-tuning of $\lambda$.
Table 5: Performance Comparison of Lagrangian Multiplier-Based Methods on Dataset PKU-SafeRLHF using Different $\lambda$ and $N$
Method $\lambda$ Average Reward Safety Rate Inference Time (s)
Alpaca-7B InferenceGuard $N=128$ 7.08 ($\pm$ 2.49) 88.14% 16.25
Best-of-N Lag. $N=128$ 0 7.92 ($\pm$ 1.43) 34.12% 8.03
1 7.42 ($\pm$ 1.72) 42.82% 7.95
2.5 6.64 ($\pm$ 1.89) 48.75% 8.89
5 5.97 ($\pm$ 1.82) 51.25% 9.62
10 5.54 ($\pm$ 1.64) 52.70% 8.88
Best-of-N Lag. $N=256$ 0 8.41 ($\pm$ 1.45) 33.86% 14.53
1 7.82 ($\pm$ 1.75) 42.95% 13.73
2.5 6.78 ($\pm$ 2.01) 51.25% 13.78
5 6.04 ($\pm$ 1.85) 54.41% 13.15
10 5.51 ($\pm$ 1.69) 55.20% 14.98
Beam Search Lag. $N=128$ 0 8.90 ($\pm$ 1.71) 27.80% 8.72
1 8.17 ($\pm$ 2.10) 41.37% 8.75
2.5 7.37 ($\pm$ 2.22) 50.46% 8.77
5 6.58 ($\pm$ 1.95) 50.19% 8.01
10 5.85 ($\pm$ 1.79) 49.93% 8.69
Beam Search Lag. + Freq. $N=128$ 1 8.15 ($\pm$ 2.09) 43.70% 8.66
2.5 7.42 ($\pm$ 2.21) 50.49% 8.74
5 6.59 ($\pm$ 1.98) 50.6% 8.42
10 5.85 ($\pm$ 1.79) 49.94% 9.55
Beam Search Lag. $N=256$ 0 9.35 ($\pm$ 1.83) 30.43% 14.31
1 8.51 ($\pm$ 2.17) 43.21% 14.63
2.5 7.58 ($\pm$ 2.25) 50.46% 14.65
5 6.69 ($\pm$ 2.08) 52.43% 15.03
10 5.90 ($\pm$ 1.82) 53.10% 14.76
Beam Search Lag. + Freq. $N=256$ 1 8.56($\pm$ 2.17) 45.22% 14.7
2.5 7.61 ($\pm$ 2.29) 50.93% 14.68
5 6.66 ($\pm$ 2.12) 52.9% 13.14
10 5.89 ($\pm$ 1.84) 52.25% 13.09
ARGS Lag. 0 6.74 ($\pm$ 1.70) 28.19% 58.58
1 4.07 ($\pm$ 1.64) 65.6% 62.28
2.5 3.98 ($\pm$ 1.61) 66.0% 69.72
5 3.21 ($\pm$ 1.59) 75.8% 63.42
10 1.23 ($\pm$ 1.63) 79.2% 61.14
Beaver-7B-v3 InferenceGuard. $N=128$ 10.26 ($\pm$ 1.42) 99.7% 9.75
Best-of-N Lag.$N=64$ 0 8.68 ($\pm$ 1.37) 77.07% 4.52
1 8.47 ($\pm$ 1.45) 81.69% 4.43
2.5 7.95 ($\pm$ 1.64) 85.11% 4.88
5 7.06 ($\pm$ 1.77) 87.48% 4.67
10 6.22 ($\pm$ 1.69) 88.14% 4.43
Best-of-N Lag.$N=128$ 0 9.15 ($\pm$ 1.32) 76.82% 7.86
1 8.92 ($\pm$ 1.43) 81.69% 8.02
2.5 8.35 ($\pm$ 1.64) 84.19% 7.72
5 7.30 ($\pm$ 1.80) 87.20% 7.92
10 6.31 ($\pm$ 1.76) 87.62% 7.82
Beam Search Lag. $N=64$ 0 11.02 ($\pm$ 1.34) 74.70% 5.18
1 10.64 ($\pm$ 1.37) 82.35% 5.9
2.5 9.99 ($\pm$ 1.58) 87.62% 5.01
5 9.84 ($\pm$ 1.4) 95.38 % 5.54
10 7.60 ($\pm$ 1.82) 89.20% 4.97
Beam Search Lag. $N=128$ 0 10.54 ($\pm$ 1.29) 74.44% 8.92
1 10.25 ($\pm$ 1.41) 82.74% 9.05
2.5 9.57 ($\pm$ 1.60) 87.10% 9.6
5 10.31 ($\pm$ 1.37) 97.36% 9.75
10 7.34 ($\pm$ 1.86) 88.41% 9.9
ARGS Lag. 0 6.72 ($\pm$ 1.83) 78.5% 67.15
1 4.01 ($\pm$ 1.61) 80.9% 75.0
2.5 3.67 ($\pm$ 1.61) 80.65% 74.5
5 2.26 ($\pm$ 1.56) 81% 73.74
10 0.95 ($\pm$ 1.67) 90.8% 68.75
Table 6: Performance Comparison of Lagrangian Multiplier-Based Methods on Datasets HEx-PHI and HH-RLHF using Different $\lambda$ and $N$ (Vicuna-7B-v1.5)
Dataset Method $\lambda$ Average Reward Safety Rate Inference Time (s)
HEx-PHI InferenceGuard. $N=128$ 6.90 ($\pm$ 2.08) 96.67% 11.01
Best-of-N Lag. $N=64$ 0 6.97 ($\pm$ 1.27) 34.33% 6.5
1 6.12 ($\pm$ 1.52) 66% 6.4
2.5 5.54 ($\pm$ 1.55) 74% 6.85
5 4.86 ($\pm$ 1.52) 78.66% 6.63
10 4.39 ($\pm$ 1.43) 80% 7.05
Best-of-N Lag. $N=128$ 0 7.83 ($\pm$ 1.19) 33% 7.92
1 6.47 ($\pm$ 1.43) 68.33% 8.46
2.5 5.84 ($\pm$ 1.51) 74.33% 9.0
5 5.22 ($\pm$ 1.39) 79.3% 9.08
10 4.62 ($\pm$ 1.37) 80.67% 8.83
Beam Search Lag. $N=64$ 0 7.98 ($\pm$ 1.62) 36% 6.8
1 6.96 ($\pm$ 1.67) 76.33% 7.29
2.5 6.28 ($\pm$ 1.66) 79.33% 7.57
5 5.36 ($\pm$ 1.42) 83% 7.34
10 4.91 ($\pm$ 1.49) 84.67% 7.11
Beam Search Lag. $N=128$ 0 8.44 ($\pm$ 1.64) 32.67% 11.83
1 7.43 ($\pm$ 1.76) 77.3% 11.8
2.5 6.53 ($\pm$ 1.75) 80.33% 11.55
5 5.70 ($\pm$ 1.57) 83% 11.28
10 5.02 ($\pm$ 1.51) 83.67% 11.41
ARGS Lag. 0 5.67 ($\pm$ 1.45) 47% 68.23
1 3.39 ($\pm$ 1.6) 83.67% 72.4
2.5 2.1 ($\pm$ 1.73) 92.67% 75.5
5 1.72 ($\pm$ 1.96) 93.33% 79.28
10 0.11 ($\pm$ 1.59) 93.33% 80.34
HH-RLHF InferenceGuard. $N=128$ 9.49 ($\pm$ 2.16) 98.97% 11.47
Best-of-N Lag. $N=64$ 0 9.14 ($\pm$ 1.99) 95.33% 5.62
1 7.80 ($\pm$ 1.89) 95.98% 6.07
2.5 7.44 ($\pm$ 2.09) 96.53% 5.71
5 6.67 ($\pm$ 2.48) 97.10% 6.83
10 5.44 ($\pm$ 2.69) 97.24% 6.64
Best-of-N Lag. $N=128$ 0 9.42 ($\pm$ 2.01) 95.6% 8.01
1 8.32 ($\pm$ 1.90) 95.84% 8.7
2.5 7.85 ($\pm$ 2.14) 96.87% 8.24
5 6.97 ($\pm$ 2.54) 97.24% 8.31
10 5.47 ($\pm$ 2.88) 97.5% 8.79
Beam Search Lag. $N=64$ 0 9.14 ($\pm$ 1.99) 95.33% 7.61
1 8.87 ($\pm$ 2.29) 96.59% 6.92
2.5 8.47 ($\pm$ 2.37) 96.96% 7.04
5 7.88 ($\pm$ 2.20) 96.6% 8.02
10 6.80 ($\pm$ 2.42) 97.51% 9.07
Beam Search Lag. + Freq. $N=64$ 1 8.91 ($\pm$ 2.34) 96.8% 6.93
2.5 8.52 ($\pm$ 2.4) 97.01% 7.04
5 7.81 ($\pm$ 2.39) 97.27% 8.54
10 6.81 ($\pm$ 2.42) 97.51% 9.01
Beam Search Lag. $N=128$ 0 9.42 ($\pm$ 2.01) 95.6% 11.4
1 9.33 ($\pm$ 2.36) 96.4% 11.06
2.5 8.89 ($\pm$ 2.51) 97.44% 11.65
5 8.05 ($\pm$ 2.25) 97.54% 11.28
10 6.85 ($\pm$ 2.56) 97.77% 12.26
Beam Search Lag. + Freq. $N=128$ 1 9.40 ($\pm$ 2.31) 96.81% 11.29
2.5 8.92 ($\pm$ 2.33) 97.96% 11.64
5 8.05 ($\pm$ 2.54) 97.64% 11.32
10 6.85 ($\pm$ 2.55) 97.71% 12.53
ARGS Lag. 0 6.83 ($\pm$ 1.83) 96.2% 70.4
1 3.98 ($\pm$ 1.79) 97% 63.78
2.5 2.65 ($\pm$ 1.8) 96.4% 67.96
5 2.02 ($\pm$ 1.79 ) 97.54% 73.74
10 0.74 ($\pm$ 1.88 ) 97.99% 74.98
Table 7: Performance Comparison of InferenceGuard w.r.t. Alpaca-7B on Dataset PKU-SafeRLHF using Different $d$, $N$, and $K$ and fixed $D=128$
Method $K$ Average Reward Average Cost Safety Rate Inference Time (s)
Alpaca-7B InferenceGuard $N=128,d=16$ 64 6.6 ($\pm$ 2.5) -0.72 94.07% 28.45
32 7.14 ($\pm$ 2.75) -0.84 94.3% 27.15
16 7.64 ($\pm$ 2.85) -0.81 94.33% 25.66
InferenceGuard $N=128,d=32$ 64 5.98 ($\pm$ 2.5) -0.86 95.65% 14.70
32 6.39 ($\pm$ 2.7) -0.94 96.3% 13.38
16 6.66 ($\pm$ 2.74) -0.89 96.05% 14.22
InferenceGuard $N=128,d=64$ 64 5.5 ($\pm$ 2.46) -0.98 96.97% 7.82
32 5.71 ($\pm$ 2.5) -0.92 96.84% 7.85
16 5.82 ($\pm$ 2.61) -0.94 96.97% 5.84
InferenceGuard $N=256,d=16$ 128 6.83 ($\pm$ 2.5) -0.88 96.18% 42.24
64 7.56 ($\pm$ 2.81) -0.98 97.1% 37.77
32 7.73 ($\pm$ 2.93) -1 98.55% 36.92
InferenceGuard $N=256,d=32$ 128 6.19 ($\pm$ 2.51) -0.99 98.15% 22.66
64 6.67 ($\pm$ 2.73) -0.94 96.97% 22.38
32 6.99 ($\pm$ 2.90) -1.03 98.15% 22.77
InferenceGuard $N=256,d=64$ 128 5.82 ($\pm$ 2.6) -0.98 98.42% 5.82
64 5.92 ($\pm$ 2.63) -1.05 99.34% 9.89
32 6.08 ($\pm$ 2.72) -1.04 97.5% 11.28
InferenceGuard $N=64,d=16$ 32 6.76 ($\pm$ 2.46) -0.5 86.56% 16.44
16 7.28 ($\pm$ 2.59) -0.65 89.2% 15.32
8 7.45 ($\pm$ 2.69) -0.6 89.06% 15.29
InferenceGuard $N=64,d=32$ 32 5.95 ($\pm$ 2.42) -0.65 90.38% 12.09
16 6.48 ($\pm$ 2.5) -0.63 90.0% 11.58
8 6.64 ($\pm$ 2.63) -0.67 90.8% 11.73
InferenceGuard $N=64,d=64$ 32 5.67 ($\pm$ 2.41) -0.73 91.17% 6.17
16 5.79 ($\pm$ 2.46) -0.76 91.57% 6.23
8 5.81 ($\pm$ 2.45) -0.75 90.6% 3.93
Robustness Analysis of InferenceGuard
To evaluate the robustness of InferenceGuard, we vary the number of samples $N$, safety budget $d$, and selection width $K$. Table 7 reports results on the PKU-SafeRLHF dataset using Alpaca-7B. We observe that increasing $N$ and reducing $d$ generally improves task reward at the expense of inference latency, while higher $d$ leads to stronger safety enforcement. For instance, at $N=256$, InferenceGuard achieves a safety rate of up to $99.34\%$ while maintaining a competitive reward of 6.08.
Figure 4: Pareto curves show the safety-reward trade-offs for decoding methods on (1) Alpaca-7B with PKU-SafeRLHF and (2) Vicuna-7B with HEx-PHI. Each curve corresponds to a $\lambda$ or safety budget ablation, tracing the approximate Pareto front.
Finally, Figure 4 shows the approximate Pareto front for reward versus safety cost across decoding strategies. InferenceGuard consistently traces a more favorable Pareto curve, offering better safety-reward trade-offs compared to baseline methods on both Alpaca-7B (PKU-SafeRLHF) and Vicuna-7B (HEx-PHI).
D.4 Limitations and Latency Analysis
While InferenceGuard demonstrates robust safety improvements across diverse models and datasets, we also acknowledge several limitations that merit further discussion.
Table 8: Win-rate Percentage Comparison on PKU-SafeRLHF evaluated by ’Deepseek-r1-distill-qwen-32b’
Method Helpfulness Win Rate (%) Harmlessness Win Rate (%)
InferenceGuard with critic 72 76.8
InferenceGuard 66.8 76.6
BeamSearch-Saute (N=256) 68.6 75
BoN-Saute (N=500) 61.4 62.0
BoN-lagrange (N=500, $\lambda=5$) 67 60.6
Args-Lagrange 14.2 52.2
RE-Control-Lagrange 52 50.4
RE-Control 50.6 49.2
ARGS-Vanilla 51.2 47.6
Figure 5: Win, tie, and loss counts of alignment methods compared against responses generated by Alpaca-7B on the PKU-SafeRLHF dataset, using Deepseek-r1-distill-qwen-32b as the judge model.
Reliance on Cost Model Quality.
Our theoretical analysis suggests that if the cost model reliably distinguishes unsafe content, InferenceGuard can almost surely generate safe responses. In practice, cost models may exhibit bias, dataset artifacts, or false positives. To address this concern, we evaluate InferenceGuard under a stronger judge model (e.g., Deepseek-r1-distill-qwen-32b[124]) independently from the cost model (as shown in Table 8 and Figure 5), and it still outperforms baseline methods in safety under distribution shift, indicating robustness and reduced overfitting to the cost function.
Table 9: Average inference time per prompt on Beaver-7B evaluated on PKU-SafeRLHF. The total inference time per prompt is decomposed into generation time using vllm and evaluation time on the reward model, cost model or critic.
Method Num Samples Beam Depth Safety Rate (%) Generation Time (s) Eval Time (s) Inference Time (Gen + Eval) (s)
Best-of-N + Augmented Safety 32 N/A 88.14% 0.82 0.6 1.42
Best-of-N + Augmented Safety 64 N/A 91.17% 1.58 1.1 2.68
Beam Search + Augmented Safety 32 64 89.60% 1.34 0.6 1.94
Beam Search + Augmented Safety 64 64 90.07% 2.67 1.2 3.87
InferenceGuard 32 64 99.21% 1.32 0.7 2.02
InferenceGuard 64 64 99.60% 2.85 1.2 4.05
Latency Overhead.
Test-time Alignment methods incur additional computation due to multi-sample decoding and online evaluation. This is an inherent challenge for most alignment methods that aim to intervene during generation. Table 9 reports end-to-end inference time across models and decoding methods. Our latency is comparable to Beam Search, particularly on fine-tuned models such as Beaver-7B-v3. We regard this gap as an “alignment tax” [27] – a computational cost required to achieve higher safety guarantees, and it is manageable based on the preference between efficiency and safety. Importantly, InferenceGuard achieves significantly better safety than methods with similar runtime profiles, and our results highlight a trade-off between latency and safety, where InferenceGuard provides strong safety guarantees at a reasonable computational cost.
D.5 Critic Network and Training Process
This section outlines the critic network architecture used for InferenceGuard. The critic network is designed to estimate the cost of partial responses and guide optimization during inference. We assume that trajectories terminate at a maximum time $T$, and the critic aims to predict the sign of the safety compliance metric $\text{z}_{T}$, and the discounted cumulative task cost $\gamma^{T}\bar{c}_{\text{task}}$.
Critic Network Architecture
The critic network takes two types of input: the hidden states ($\bm{h}_{t}$) and the key-value pairs ($\mathbf{o}_{t}$), representing contextual and state information, respectively. These are passed through a series of layers to estimate the required outputs. The network utilizes downscaling and attention layers to reduce the dimensionality of the input data, ensuring efficient processing of large-scale representations.
In terms of model size, the total parameter count of the critic network is approximately 0.7 billion parameters, providing a balance between model capacity and computational efficiency.
Hyperparameter Value
Hidden Dimension 4096
Learning Rate $1\times 10^{-5}$
Number of Epochs 50
Discount Factor ($\gamma$) 0.999
Batch Size 8
Safety Budget $d$ 10
Table 10: Hyperparameters for Critic Network Training.
Training Process
The critic network is trained using a combination of optimization techniques aimed at predicting safety compliance and task cost accurately. The network is optimized with the hyperparameters in Table 10.
During training, the network is fed batches of hidden states and key-value pairs, and the weights are updated to minimize the loss between predicted and true values. The critic network’s ability to predict both the safety compliance and task cost ensures it can guide the optimization process at inference time, while adhering to safety constraints.
The model uses a penalty term to enforce the safety budget constraint. This penalty discourages the network from violating the safety threshold, steering for safer responses during intervention.
D.6 Qualitative Comparisons
We present several examples from the PKU-SafeRLHF test dataset to demonstrate how InferenceGuard steers the base model outputs towards safer responses. These examples highlight differences in response safety between InferenceGuard and baseline methods, including BoN, Beam Search, RE-Control, and ARGS, evaluated on the Alpaca and Beaver-v3 models, as shown in Figure 6 and 7. In each case, InferenceGuard successfully adheres to safety constraints while maintaining task performance.
(a) Alpaca Example 1
(b) Alpaca Example 2
Figure 6: Generated response by different methods on the Alpaca-7B
(a) Beaver Example 1
(b) Beaver Example 2
Figure 7: Generated response by different methods on the Beaver-v3-7B
Appendix E Broader Impact Statement
This work contributes to the safe and responsible deployment of large language models (LLMs) by developing InferenceGuard, an inference-time alignment method that ensures almost surely safe responses. Given the increasing reliance on LLMs across various domains, including healthcare, education, legal systems, and autonomous decision-making, guaranteeing safe and aligned outputs is crucial for mitigating misinformation, bias, and harmful content risks.
To further illustrate the effectiveness of our approach, we have included additional examples in the appendix demonstrating that our method successfully produces safe responses. These examples were generated using standard prompting with available large language models LLMs. Additionally, we have added a warning at the beginning of the manuscript to inform readers about the nature of these examples. Our primary motivation for this addition is to highlight the safety improvements achieved by our method compared to existing alternatives. We do not foresee these examples being misused in any unethical manner, as they solely showcase our model’s advantages in ensuring safer AI interactions. Finally, we emphasize that our method is designed specifically to enhance AI safety, and as such, we do not anticipate any potential for unethical applications.
InferenceGuard enhances the scalability and adaptability of safe AI systems by introducing a formally grounded safety mechanism that does not require model retraining while reducing the resource costs associated with traditional RLHF methods. The proposed framework advances AI safety research by providing provable safety guarantees at inference time, an area that has received limited attention in prior work.
While this method significantly improves safety in LLM outputs, it does not eliminate all potential risks, such as adversarial manipulation or emergent biases in model responses. Future work should explore robustness to adversarial attacks, contextual fairness, and ethical considerations in deploying safety-aligned LLMs across different cultural and regulatory landscapes. Additionally, transparency and accountability in AI safety mechanisms remain essential for gaining public trust and ensuring alignment with societal values. This work aims to empower developers and policymakers with tools for ensuring safer AI deployment while contributing to the broader conversation on AI ethics and governance.